### v1

# Vanishing criteria for Ceresa cycles

Abstract.
Let $C$ be a smooth projective curve, and let $J$ be its Jacobian. We prove vanishing criteria for the Ceresa cycle $\kappa(C)\in\mathrm{CH}_{1}(J)\otimes{\rm\mathbb{Q}}$ in the Chow group of 1-cycles on $J$. Namely,
  1. $(A)$
    If $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$, then $\kappa(C)$ vanishes;
  2. $(B)$
    If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then $\kappa(C)$ vanishes modulo algebraic equivalence,
with criterion $(B)$ conditional on the Hodge conjecture.
We then study the first interesting case where $(B)$ holds but $(A)$ does not, namely the case of Picard curves $C\colon y^{3}=x^{4}+ax^{2}+bx+c$. Using work of Schoen on the Hodge conjecture, we show that the Ceresa cycle of a Picard curve is torsion in the Griffiths group. Moreover, we determine exactly when it is torsion in the Chow group. As a byproduct, we show that there are infinitely many plane quartic curves over ${\rm\mathbb{Q}}$ with torsion Ceresa cycle (in fact, there is a one parameter family of such curves). Finally, we determine which automorphism group strata are contained in the vanishing locus of the universal Ceresa cycle over $\mathcal{M}_{3}$.
1. Introduction
Let $k$ be an algebraically closed field and $C$ a smooth, projective, and connected curve over $k$ of genus $g\geq 2$ with Jacobian variety $J$. Let $e$ be a degree-$1$ divisor of $C$ and let $\iota_{e}\colon C\hookrightarrow J$ be the Abel-Jacobi map based at $e$. We study the torsion behaviour of the Ceresa cycle
(1.1) $\kappa_{C,e}:=[\iota_{e}(C)]-(-1)^{*}[\iota_{e}(C)]\in{\rm CH}_{1}(J)$
in the Chow group modulo rational equivalence. If $\kappa_{C,e}$ is torsion, then $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$, where $K_{C}$ is the canonical divisor class. Moreover if $(2g-2)e$ is canonical, then the image of $\kappa_{C,e}$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$ is independent of $e$ and we denote it by $\kappa(C)$, see §2.7 for these claims. The class $\kappa(C)$ vanishes if and only if $\kappa_{C,e}$ is torsion for some degree-$1$ divisor $e$.
We also consider the image $\bar{\kappa}(C)$ of $\kappa(C)$ in the Griffiths group ${\rm Gr}_{1}(J)\otimes{\rm\mathbb{Q}}$ of homologically trivial $1$-cycles modulo algebraic equivalence. When $g=2$, or more generally when $C$ is hyperelliptic, it is easy to see that $\kappa(C)=0$. On the other hand, Ceresa famously showed that $\bar{\kappa}(C)\neq 0$ for a very general curve $C$ over ${\rm\mathbb{C}}$ of genus $g\geq 3$ [14].
The vanishing of the Ceresa cycle is interesting for various reasons. For example, $\kappa(C)=0$ if and only if the Chow motive $\mathfrak{h}(C)$ has a multiplicative Chow-Künneth decomposition (by [19, Proposition 3.1] and Proposition 2.12). Moreover $\bar{\kappa}(C)=0$ if and only if the tautological subring modulo algebraic equivalence is generated by a theta divisor (by [6, Corollary 3.4]), in which case Poincaré’s formula $[C]=\frac{\Theta^{g-1}}{(g-1)!}$ holds modulo algebraic equivalence. More generally, the Ceresa cycle over $\mathcal{M}_{g}$ serves as a testing ground for the study of homologically trivial algebraic cycles in codimension greater than 1.
1.1. Vanishing criteria
We prove cohomological vanishing criteria for Ceresa cycles of curves with nontrivial automorphisms. Let $\operatorname{H}^{*}(-)$ be a Weil cohomology functor, such as $\ell$-adic cohomology with $\ell\neq{\rm char\hskip 2.84526pt}(k)$ or singular cohomology when $k={\rm\mathbb{C}}$. Note that the finite group ${\rm Aut}(C)$ acts on $\operatorname{H}^{*}(J)$, by functoriality. Cupping with the principal polarization gives an injection $\operatorname{H}^{1}(J)(-1)\hookrightarrow\operatorname{H}^{3}(J)$, allowing us to define the primitive cohomology $\operatorname{H}^{3}(J)_{\mathrm{prim}}:=\operatorname{H}^{3}(J)/\operatorname{H}^{1}(J)(-1)$.
Theorem A.
If $\operatorname{H}^{3}(J)_{\mathrm{prim}}^{{\rm Aut}(C)}=0$, then $\kappa(C)=0$.
This improves on a recent result of Qiu and Zhang stating that if $(\operatorname{H}^{1}(C)^{\otimes 3})^{{\rm Aut}(C)}=0$, then $\kappa(C)=0$ [39].111 This formulation is equivalent to theirs, using [46, Theorem 1.5.5]. By contrast, Theorem A requires only the weaker condition that the subrepresentation $\operatorname{H}^{3}(J)_{\mathrm{prim}}\subset\operatorname{H}^{3}(J)\simeq\bigwedge^{3}\operatorname{H}^{1}(C)\subset\operatorname{H}^{1}(C)^{\otimes 3}$ has no nontrivial ${\rm Aut}(C)$-fixed points. If the quotient $C/{\rm Aut}(C)$ has genus 00, then our hypothesis is equivalent to $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$.
Our proof of Theorem A is inspired by Beauville’s proof that for the curve $y^{3}=x^{4}+x$, the image of $\kappa(C)$ under the complex Abel-Jacobi map vanishes [7]. To achieve a vanishing result in the Chow group, we work directly with the rational Chow motive $\mathfrak{h}^{3}(J)$ and make crucial use of the finite-dimensionality results of Kimura [25].
Our second result is a vanishing criterion for $\overline{\kappa}(C)$, however our proof works only in characteristic 00 and is conditional on the Hodge conjecture:
Theorem B.
Suppose $\mathrm{char}(k)=0$ and assume the Hodge conjecture for abelian varieties. If $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, then $\bar{\kappa}(C)=0$.
More precisely, we require the Hodge conjecture for $J\times A$, for a specific abelian variety $A$ described in Proposition 3.3. Note that $\operatorname{H}^{0}(J,\Omega_{J}^{3})\simeq\bigwedge^{3}\operatorname{H}^{0}(C,\Omega_{C}^{1})$, so the conditions of Theorems A (when $\mathrm{char}(k)=0$) and B both depend only on the abstract representation $(G,V)=({\rm Aut}(C),\operatorname{H}^{0}(C,\Omega_{C}^{1}))$.
The proof of Theorem B is in the same spirit as that of Theorem A. The Hodge conjecture is used to show that the motive $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}$ is isomorphic to $\mathfrak{h}^{1}(A)(-1)$, from which the algebraic triviality of $\kappa(C)$ follows.
1.2. Picard curves
The condition $\operatorname{H}^{3}(J)_{\mathrm{prim}}^{{\rm Aut}(C)}=0$ in Theorem A is only rarely satisfied, e.g. it holds for exactly two plane quartic curves over ${\rm\mathbb{C}}$ (see Theorem D below). However, the condition $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$ of Theorem B is satisfied for all Picard curves $y^{3}=x^{4}+ax^{2}+bx+c$. For such curves it turns out that we require a nontrivial case of the Hodge conjecture to apply Theorem B, namely that a certain Weil class on an abelian fourfold is algebraic. Fortunately, Schoen has proved the Hodge conjecture in our specific situation [41]. (See also recent work of Markman [32].) This leads to an unconditional proof of the vanishing of $\bar{\kappa}(C)$. By further analyzing $\kappa(C)$ we prove the following precise characterization of the vanishing of $\kappa(C)$ in the Picard family:
Theorem C.
Suppose $\mathrm{char}(k)=0$ and let $C_{f}$ be a smooth projective Picard curve with model
(1.2) $C_{f}\colon y^{3}=f(x)=x^{4}+ax^{2}+bx+c$
for some $a,b,c\in k$. Consider the point $P_{f}=(a^{2}+12c,72ac-2a^{3}-27b^{2})$ on the elliptic curve
$E_{f}\colon y^{2}=4x^{3}-27\cdot{\rm disc}(f).$
Then $\bar{\kappa}(C_{f})=0$, and $\kappa(C_{f})=0$ if and only if $P_{f}\in E_{f}(k)$ is torsion.
Here, ${\rm disc}(f)$ is the discriminant of $f$ and the coordinates of $P_{f}$ are the usual $I$- and $J$-invariants of the binary quartic form $z^{4}f(x/z)$; see §4.1. The fact that $P_{f}$ defines a point on $E_{f}$ follows from a classical relation between the invariants of a binary quartic. See also §4.2 for a more precise version of Theorem C, which relates order of $\kappa_{C_{f},P_{\infty}}$ to the order of $P_{f}$, and in particular shows that $\kappa_{C_{f},P_{\infty}}$ can have arbitrarily large (finite) order in this family. Theorem C also gives one-parameter families of plane quartic curves with vanishing $\kappa(C)$, corresponding to components of the torsion locus of $P_{f}$. Such families exist even over ${\rm\mathbb{Q}}$, e.g. the family $y^{3}=x^{4}-12x^{2}+tx-12$.
Theorem C generalizes the results of [27] concerning bielliptic Picard curves, i.e. those with $b=0$ in (1.2). There we exploited the explicit geometry of the bielliptic cover to characterize the vanishing locus of $\kappa(C)$. Here we must proceed more indirectly: using Schoen’s work on the Hodge conjecture, we first show that the vanishing of $\kappa(C)$ is equivalent to the vanishing of its Abel–Jacobi image in the intermediate Jacobian. Using some formal considerations concerning isotrivial families of elliptic curves, we are able to bootstrap the calculations of [27] to determine the Abel-Jacobi image up to multiple and obtain the characterization of Theorem C.
In §3, we give two examples of genus 4 families of curves $\mathcal{C}\to B$ for which Theorem B applies. It would be interesting to completely characterize the vanishing of $\kappa(C)$ in these families as well, and to find more such families. The general proof strategy would follow that of Theorem C: using the Hodge conjecture, one shows that the vanishing of $\kappa(\mathcal{C}_{b})$ is equivalent to the vanishing of $\sigma(b)$, where $\sigma$ is a section of a certain abelian scheme $\mathcal{A}\to B$. (The analytification of $\sigma$ is the normal function associated to the Ceresa cycle over $B$, which lands in the ${\rm Aut}(C)$-fixed points of the intermediate Jacobian.) One then tries to identify $\mathcal{A}\to B$ and $\sigma$ explicitly.
1.3. Vanishing loci in genus $3$
Let $\mathcal{M}_{g}$ be the coarse moduli space of genus-$g$ curves in characteristic zero, seen as a variety over ${\rm\mathbb{Q}}$. Let $V_{g}^{\mathrm{rat}}\subset\mathcal{M}_{g}$ be the subset consisting of all (geometric) isomorphism classes of curves with vanishing $\kappa(C)$. Let $V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ be the analogous subset for $\bar{\kappa}(C)$. The sets $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ are both countable unions of closed algebraic subvarieties (Lemma 5.1). Moreover $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}$ and Ceresa’s result [14] shows that $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ for all $g\geq 3$. What can be said about the irreducible components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$ and their dimensions? Collino and Pirola have shown that $V_{3}^{\mathrm{rat}}$ does not contain subvarieties of dimension $\geq 4$ that are not themselves contained in the hyperelliptic locus [16, Corollary 4.3.4]. Theorem C shows that $V_{3}^{\mathrm{alg}}$ contains the $2$-dimensional Picard locus and that $V_{3}^{\mathrm{rat}}$ contains infinitely many $1$-dimensional components.
Recall that there is a stratification of the non-hyperelliptic locus of $\mathcal{M}_{3}$ by locally closed subvarieties $X_{G}$, indexed by certain finite groups $G$, with the property that a non-hyperelliptic curve $[C]\in\mathcal{M}_{3}({\rm\mathbb{C}})$ lies in $X_{G}$ if and only if ${\rm Aut}(C)\simeq G$. It turns out that the isomorphism class of the ${\rm Aut}(C)$-representation $\operatorname{H}^{0}(C,\Omega_{C}^{1})$ does not depend on the choice of curve $[C]$ in $X_{G}({\rm\mathbb{C}})$. Our final theorem determines exactly which $X_{G}$ are contained in $V_{3}^{\mathrm{rat}}$ or $V_{3}^{\mathrm{alg}}$.
Theorem D.
Let $G$ be a finite group isomorphic to the automorphism group of a plane quartic. Then $X_{G}\subset V_{3}^{\mathrm{rat}}$ if and only if $G=C_{9}$ or $G_{48}$, and $X_{G}\subset V_{3}^{\mathrm{alg}}$ if and only if $X_{G}$ is contained in the Picard locus, in other words if and only if $G=C_{3},C_{6},C_{9}$ or $G_{48}$.
Here, $G_{48}$ is the group with GAP label $(48,33)$. The strata $X_{C_{9}}$ and $X_{G_{48}}$ are single closed points, represented by the curves $y^{3}=x^{4}-x$ and $y^{3}=x^{4}+1$ respectively.
Theorem D shows that $X_{G}\subset V_{3}^{\mathrm{rat}}$ if and only if $\operatorname{H}^{3}(J)^{G}=0$ for every (equivalently, some) curve in $X_{G}$, and that $X_{G}\subset V_{3}^{\mathrm{alg}}$ if and only if $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{G}=0$ for every curve in $X_{G}$. Thus, for $g=3$, the criteria of Theorems A and B exactly single out the strata $X_{G}$ where $\kappa(C)$ or $\bar{\kappa}(C)$ identically vanishes.
Finally, Shou-wu Zhang has recently announced a proof of a Northcott property for the Beilinson-Bloch height of the modified diagonal cycle $\Delta_{GKS}$ on $C^{3}$. More precisely, for each $g\geq 3$, there exists an open dense subset $U_{g}\subset\mathcal{M}_{g}$ such that for any $X\in{\rm\mathbb{R}}$ and $d\in{\rm\mathbb{N}}$, the number of $C\in U_{g}(\bar{\rm\mathbb{Q}})$ defined over a number field of degree at most $d$ and with $\langle\Delta_{GKS},\Delta_{GKS}\rangle<X$ is finite. In order to better understand Ceresa vanishing loci in families of curves, it would be of great interest to try to identify the largest such open dense set $U_{g}$, or equivalently, its complement $Z_{g}:=\mathcal{M}_{g}\setminus U_{g}$. The Northcott property implies that any positive dimensional component of $V_{g}^{\mathrm{rat}}$ is contained in $Z_{g}$, but there may be other components of $Z_{g}$ as well. Indeed, Theorem C shows that $X_{C_{3}}\subset Z_{3}$, even though $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$; see Remark 5.4.
1.4. Structure of paper
In §2 we collect some standard results on Chow groups, Chow motives and Ceresa cycles. In §3 we prove the cohomological vanishing criteria (Theorems A and B) and give some examples. In §4 we study the family of Picard curves in detail and prove Theorem C. Finally, in §5 we introduce the Ceresa vanishing loci $V_{g}^{\mathrm{rat}},V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ and determine which automorphism strata they contain in genus $3$, proving Theorem D.
1.5. Acknowledgements
We thank Jeffrey Achter, Ben Moonen, Burt Totaro and Congling Qiu for helpful conversations. This research was carried out while the first author was a Research Fellow at St John’s College, University of Cambridge, which he thanks for providing excellent working conditions. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157).
2. Notation and background
2.1. Chow groups
Let $k$ be field. A variety is by definition a separated scheme of finite type over $k$. We say a variety is nice if it is smooth, projective and geometrically integral. If $X$ is a smooth and geometrically integral variety and $p\in\{0,\dots,\dim X\}$, let ${\rm CH}^{p}(X)$ denote the Chow group (with ${\rm\mathbb{Z}}$-coefficients) of codimension $p$ cycles modulo rational equivalence. If $Z\subset X$ is a closed subscheme of codimension $p$, we denote its class in ${\rm CH}^{p}(X)$ by $[Z]$ (using [44, Tag 02QS] if $Z$ is not integral).
If $X$ is additionally projective, then ${\rm CH}^{p}(X)$ has a filtration by subgroups
${\rm CH}^{p}(X)_{\mathrm{alg}}\subset{\rm CH}^{p}(X)_{\mathrm{hom}}\subset{\rm CH}^{p}(X),$
where ${\rm CH}^{p}(X)_{\mathrm{alg}}$ is the subgroup of algebraically trivial cycles (in the sense of [2, §3.1]) and ${\rm CH}^{p}(X)_{\mathrm{hom}}$ the subgroup of homologically trivial cycles (with respect to a fixed Weil cohomology theory for nice varieties over $k$). The Griffiths group is by definition ${\rm Gr}^{p}(X)={\rm CH}^{p}(X)_{\mathrm{hom}}/{\rm CH}^{p}(X)_{\mathrm{alg}}$. We occasionally write ${\rm CH}_{p}(X)={\rm CH}^{\dim X-p}(X)$ and ${\rm Gr}_{p}(X)={\rm Gr}^{\dim X-p}(X)$. If $R$ is a ring, we write ${\rm CH}^{p}(X)_{R}={\rm CH}^{p}(X)\otimes_{{\rm\mathbb{Z}}}R$ and ${\rm Gr}^{p}(X)_{R}={\rm Gr}^{p}(X)\otimes_{{\rm\mathbb{Z}}}R$.
2.2. Base change, specialization and families
We state three lemmas concerning operations on cycles. These seem to be standard, but we could not locate proofs in the literature.
Lemma 2.1.
Let $X/k$ be a nice variety and $K/k$ a (not necessarily finite) extension of fields.
  1. (1)
    The base change maps ${\rm CH}^{p}(X)_{\rm\mathbb{Q}}\rightarrow{\rm CH}^{p}(X_{K})_{\rm\mathbb{Q}}$ and ${\rm Gr}^{p}(X)_{\rm\mathbb{Q}}\rightarrow{\rm Gr}^{p}(X_{K})_{\rm\mathbb{Q}}$ are injective.
  2. (2)
    If in addition $k$ is algebraically closed, the base change maps ${\rm CH}^{p}(X)\rightarrow{\rm CH}^{p}(X_{K})$ and ${\rm Gr}^{p}(X)\rightarrow{\rm Gr}^{p}(X_{K})$ are injective.
Proof.
The proof is an adaptation of [12, Lemma 1A.3, p. 22]. We first prove $(2)$ for ${\rm CH}^{p}(X)\rightarrow{\rm CH}^{p}(X_{K})$. Suppose $\alpha\in{\rm CH}^{p}(X)$ has trivial image in ${\rm CH}^{p}(X_{K})$. Then $\alpha$ already has trivial image in ${\rm CH}^{p}(X_{K^{\prime}})$, where $K^{\prime}\subset K$ is a subfield that is finitely generated over $k$, since the data witnessing triviality in ${\rm CH}^{p}(X_{K})$ can be defined over such a subfield. By spreading out, we can find a smooth integral variety $U/k$ with function field $K^{\prime}$ such that $\alpha$ has trivial image in ${\rm CH}^{p}(X\times_{k}U)$. Since $k$ is algebraically closed, there exists a $k$-point $u\in U(k)$. Pulling back along $u$ defines a left-inverse ${\rm CH}^{p}(X\times_{k}U)\rightarrow{\rm CH}^{p}(X)$ to the map ${\rm CH}^{p}(X)\rightarrow{\rm CH}^{p}(X\times_{k}U)$. It follows that $\alpha$ is trivial in ${\rm CH}^{p}(X)$, as desired. The argument for ${\rm Gr}^{p}(X)$ is identical and omitted.
We now prove $(1)$ for ${\rm CH}^{p}(X)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{p}(X_{K})_{{\rm\mathbb{Q}}}$. There exists a field $L$ containing both $K$ and an algebraic closure $\bar{k}$ of $k$. It therefore suffices to prove the two base change maps ${\rm CH}^{p}(X)_{\rm\mathbb{Q}}\rightarrow{\rm CH}^{p}(X_{\bar{k}})_{\rm\mathbb{Q}}\rightarrow{\rm CH}^{p}(X_{L})_{{\rm\mathbb{Q}}}$ are both injective. The first one follows from the fact that for a finite extension $k^{\prime}/k$ the pushforward map ${\rm CH}^{p}(X_{k^{\prime}})\rightarrow{\rm CH}^{p}(X)$, when precomposed with the base change map, is multiplication by $[k^{\prime}:k]$. The second follows from Part $(2)$. The case of ${\rm Gr}^{p}(X)_{{\rm\mathbb{Q}}}$ is again analogous. ∎
To discuss specialization in the next lemma, let $R$ be a discrete valuation ring with fraction field $K$ and residue field $k$. Let $X\rightarrow{\rm Spec}(R)$ be a smooth, projective morphism with geometrically integral fibers, so the generic and special fibers $X_{K}$ and $X_{k}$ are nice varieties over $K$ and $k$ respectively. In this setting, Fulton has defined [20, §20.3] a specialization morphism $\mathrm{sp}\colon{\rm CH}^{p}(X_{K})\rightarrow{\rm CH}^{p}(X_{k})$ for every $0\leq p\leq\dim(X_{K})$. It has the property that if $Z\subset X$ is a closed integral subscheme of codimension $p$, flat over $R$, then $\mathrm{sp}([Z_{K}])=[Z_{k}]$.
Lemma 2.2.
In the above notation, $\mathrm{sp}$ sends ${\rm CH}^{p}(X_{K})_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ to ${\rm CH}^{p}(X_{k})_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$.
Proof.
Let $C/K$ be a nice curve with $K$-points $t_{0},t_{1}\in C(K)$ and let $Z\subset(X_{K})\times_{K}C$ be an integral closed subscheme of codimension $p$, flat over $C$. By [2, Theorem 1] and the remarks thereafter, it suffices to prove that $\mathrm{sp}([Z_{t_{0}}]-[Z_{t_{1}}])\in{\rm CH}^{p}(X_{k})_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ for each such tuple $(C,t_{0},t_{1},Z)$.
Let $\mathcal{C}\rightarrow{\rm Spec}(R)$ be an integral, projective, flat, proper and regular model of $C$, which exists by [30, Proposition 10.1.8]. Let $\mathrm{Hilb}_{X/R}\rightarrow{\rm Spec}(R)$ be the Hilbert scheme of the projective morphism $X\rightarrow{\rm Spec}(R)$. The closed subscheme $Z$ determines a $K$-morphism $\alpha\colon C\rightarrow\mathrm{Hilb}_{X/R}$. Since $C$ is connected, the image of $\alpha$ is contained in an open and closed subscheme $\mathrm{Hilb}_{X/R}^{\phi}$ with fixed Hilbert polynomial. The surface $\mathcal{C}$ is regular, $\mathrm{Hilb}^{\phi}_{X/R}\rightarrow{\rm Spec}(R)$ is projective [3, Corollary (2.8)] and we may view $\alpha$ as a rational map $\mathcal{C}\dashrightarrow\mathrm{Hilb}_{X/R}^{\phi}$. Therefore by [30, Theorem 9.2.7] there exists a birational map $\widetilde{\mathcal{C}}\rightarrow\mathcal{C}$, obtained by successively blowing up closed points in the special fiber, and an $R$-morphism $\tilde{\alpha}\colon\widetilde{\mathcal{C}}\rightarrow\mathrm{Hilb}^{\phi}_{X/R}$ with generic fiber $\alpha$. In other words, there exists a closed subscheme $\mathcal{Z}\subset\widetilde{\mathcal{C}}\times_{R}X$, flat over $\widetilde{\mathcal{C}}$, whose generic fiber equals $Z$.
The $K$-points $t_{i}\in C(K)$ extend to $R$-points $\tilde{t}_{i}\colon{\rm Spec}(R)\rightarrow\widetilde{\mathcal{C}}$ whose reductions $\bar{t}_{i}\in\widetilde{\mathcal{C}}_{k}(k)$ land in the smooth locus of $\widetilde{\mathcal{C}}_{k}$. The closed subschemes $\mathcal{Z}_{\tilde{t}_{i}}\subset X$ are flat over $R$. Therefore $\mathrm{sp}([Z_{t_{0}}]-[Z_{t_{1}}])=[\mathcal{Z}_{\bar{t}_{0}}]-[\mathcal{Z}_{\bar{t}_{1}}]$.
By the Zariski connectedness theorem, the special fiber $\widetilde{\mathcal{C}}_{k}$ is connected. Consequently, by resolving irreducible components of $\widetilde{\mathcal{C}}_{k}$, we can connect $\bar{t}_{0}$ to $\bar{t}_{1}$ by a sequence of nice curves: there exists a finite extension $k^{\prime}/k$, a collection of nice curves $D_{1},\dots,D_{n}$ over $k^{\prime}$, and for each $1\leq i\leq n$ a pair of points $s_{i,1},s_{i,2}\in D_{i}(k^{\prime})$ and a morphism $\varphi_{i}\colon D_{i}\rightarrow\mathcal{C}_{k^{\prime}}$ such that $\varphi_{1}(s_{1,1})=\bar{t}_{1}$, $\varphi_{i}(s_{i,2})=\varphi_{i+1}(s_{i+1,1})$ for all $11\leq i\leq n-$ 1 and $\varphi_{n}(s_{n,2})=\bar{t}_{2}$. Letting $W^{(i)}$ be the pullback of $\mathcal{Z}_{k}$ along $\varphi_{i}$, we have
$\mathrm{sp}([Z_{t_{0}}]-[Z_{t_{1}}])_{k^{\prime}}=([W^{(1)}_{s_{1,1}}]-[W^{(1)}_{s_{1,2}}])+\cdots+([W^{(n)}_{s_{n,1}}]-[W^{(n)}_{s_{n,2}}]).$
Since each term $([W^{(i)}_{s_{i,1}}]-[W^{(i)}_{s_{i,2}}])$ lies in ${\rm CH}^{p}(X_{k^{\prime}})_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ by definition of algebraic triviality, the same is true for their sum. Taking the pushforward along $X_{k^{\prime}}\rightarrow X_{k}$, we see that $[k^{\prime}:k]\cdot\mathrm{sp}([Z_{t_{0}}]-[Z_{t_{1}}])\in{\rm CH}^{p}(X_{k})_{\mathrm{alg}}$, as desired. ∎
Remark 2.3.
The proof shows that if $K$ is perfect and $k$ is algebraically closed, then $\mathrm{sp}$ even sends ${\rm CH}^{p}(X_{K})_{\mathrm{alg}}$ to ${\rm CH}^{p}(X_{k})_{\mathrm{alg}}$.
Lemma 2.4.
Let $X\rightarrow S$ be a smooth proper morphism of smooth varieties over a field $k$. Let $\alpha$ be a codimension $p$ cycle on $X$. Then the locus of points $s\in S$ such that the (Gysin) fiber $\alpha_{s}\in{\rm CH}^{p}(X_{s})_{{\rm\mathbb{Q}}}$ is zero (resp. lies in ${\rm CH}^{p}(X_{s})_{\mathrm{alg},{\rm\mathbb{Q}}}$) is a countable union of closed algebraic subvarieties of $S$.
Proof.
There exists a countable subfield $k_{0}\subset k$, a smooth proper morphism $X_{0}\rightarrow S_{0}$ of varieties over $k_{0}$ and a cycle $\alpha_{0}$ on $X_{0}$ whose base change to $k$ are $X\rightarrow S$ and $\alpha$ respectively. Since $k_{0}$ is countable, $S_{0}$ has only countably many closed subschemes. Let $\mathcal{F}$ be the collection of integral closed subschemes $Z\subset S_{0}$ such that $\alpha_{0}$ is zero in ${\rm CH}^{p}(X_{0,\eta(Z)})_{{\rm\mathbb{Q}}}$, where $\eta(Z)\in S_{0}$ denotes the generic point of $Z$. Then we claim that the locus of $s\in S$ for which $\alpha_{s}\in{\rm CH}^{p}(X_{s})_{{\rm\mathbb{Q}}}$ is zero is exactly the union $\bigcup_{Z\in\mathcal{F}}Z_{k}\subset S$. This follows from Lemmas 2.1 and properties of the specialization map; we omit the details, and the similar argument for algebraic triviality. ∎
2.3. Chow motives
We recall a few relevant facts about the category $\mathsf{Mot}(k)$ of (pure, contravariant) Chow motives $M$ over $k$ with ${\rm\mathbb{Q}}$-coefficients; see [42] or [38, §2] for basic definitions.
Denote the Lefschetz motive by $\mathbb{L}$, its $n$th tensor power by $\mathbb{L}^{n}$ and write $M(n)=M\otimes\mathbb{L}^{n}$. Following [38, §2.5], the Chow group in codimension $p$ of $M$ is by definition ${\rm CH}^{p}(M)={\rm Hom}_{\mathsf{Mot}(k)}(\mathbb{L}^{p},M)$. If $M=\mathfrak{h}(X)$ where $X$ is smooth projective over $k$, then ${\rm CH}^{p}(M)={\rm CH}^{p}(X)_{{\rm\mathbb{Q}}}$. (Beware that we need to take ${\rm\mathbb{Q}}$-coefficients on the right hand side.) If $C$ is a nice curve over $k$ and $\varphi\colon\mathfrak{h}(C)\rightarrow M(p-1)$ a morphism in $\mathsf{Mot}(k)$, we obtain a homomorphism of abelian groups ${\rm CH}^{1}(\varphi)\colon{\rm CH}^{1}(C)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{p}(M)$. We define ${\rm CH}^{p}(M)_{\mathrm{alg}}$ to be the union of the images of ${\rm CH}^{1}(\varphi)$ ranging over all pairs $(C,\varphi)$ as above. This coincides with ${\rm CH}^{p}(X)_{\mathrm{alg}}\otimes{\rm\mathbb{Q}}$ when $M=\mathfrak{h}(X)$, see [2, Corollary 3.13].
Finally, we define motives of fixed points. If $G$ is a finite group acting on a motive $M$, write $M^{G}$ for the submotive cut out by the idempotent $\frac{1}{\#G}\sum_{g\in G}g_{*}\in{\rm End}(M)$. We have ${\rm CH}^{p}(M^{G})={\rm CH}^{p}(M)^{G}$. If $G$ acts on a nice variety $X$, then $G$ acts on $\mathfrak{h}(X)$ and ${\rm CH}^{p}(X)$, and ${\rm CH}^{p}(\mathfrak{h}(X)^{G})={\rm CH}^{p}(X)^{G}\otimes{{\rm\mathbb{Q}}}$.
2.4. Chow–Künneth decomposition for abelian varieties
Let $A/k$ be a $g$-dimensional abelian variety. Deninger and Murre [18, §3] have constructed a canonical Chow–Künneth decomposition
(2.1) $\mathfrak{h}(A)=\bigoplus_{i=0}^{2g}\mathfrak{h}^{i}(A),$
uniquely characterized by the following property: if $(n)\colon A\rightarrow A$ denotes the multiplication-by-$n$, then $(n)^{*}$ acts on $\mathfrak{h}^{i}(A)$ via $n^{i}$ for every integer $n$. On the other hand, Beauville [5] has shown that there exists a direct sum decomposition ${\rm CH}^{p}(A)_{\rm\mathbb{Q}}=\bigoplus_{s=p-g}^{p}{\rm CH}^{p}_{(s)}(A)$, where
(2.2) ${\rm CH}^{p}_{(s)}(A)=\{\alpha\in{\rm CH}^{p}(A)_{\rm\mathbb{Q}}\colon(n)^{*}\alpha=n^{2p-s}\alpha\quad\forall n\in{\rm\mathbb{Z}}\}.$
The two decompositions are linked by the formula ${\rm CH}^{p}(\mathfrak{h}^{i}(A))={\rm CH}^{p}_{(2p-i)}(A)$. Beauville conjectured that ${\rm CH}^{p}_{(s)}(A)=0$ when $s<0$, and he proved it when $p\in\{0,1,g-2,g-1,g\}$ [5, Proposition 3(a)].
Example 2.5.
For $p=1$, the Beauville decomposition ${\rm CH}^{1}(A)_{\rm\mathbb{Q}}={\rm CH}^{1}_{(0)}(A)\oplus{\rm CH}^{1}_{(1)}(A)$ is the decomposition of a divisor class into symmetric and anti-symmetric classes.
Lemma 2.6.
If $A/k$ is an abelian variety, then ${\rm CH}^{p}(\mathfrak{h}^{1}(A))_{\mathrm{alg}}={\rm CH}^{p}(\mathfrak{h}^{1}(A))$ for all $p\geq 0$.
Proof.
By Lemma 2.1(1), we may assume $k$ is algebraically closed. The only nonzero Chow group of $\mathfrak{h}^{1}(A)$ is ${\rm CH}^{1}(\mathfrak{h}^{1}(A))={\rm CH}^{1}_{(1)}(A)$, the set of anti-symmetric elements of ${\rm CH}^{1}(A)_{{\rm\mathbb{Q}}}$. The lemma follows from the fact that ${\rm CH}^{1}_{(1)}(A)={\rm CH}^{1}(A)_{\mathrm{hom}}\otimes{\rm\mathbb{Q}}$ and that homological and algebraic equivalence coincide for codimension-$1$ cycles. ∎
2.5. The Lefschetz decomposition for abelian varieties
Let $A/k$ be an abelian variety with polarization $\lambda\colon A\rightarrow A^{\vee}$. Let $\ell\in{\rm CH}^{1}_{(0)}(A)$ be the unique class satisfying $2\ell=(1,\lambda)^{*}\mathcal{P}$, where $\mathcal{P}\in{\rm Pic}(A\times A^{\vee})$ is the Poincaré bundle. Künnemann has shown [26, Theorem 5.2] that intersecting with $\ell^{g-i}$ induces an isomorphism $\mathfrak{h}^{i}(A)\rightarrow\mathfrak{h}^{2g-i}(A)(g-i)$ for all $0\leq i\leq g$. This induces a Lefschetz decomposition of the Chow–Künneth components $\mathfrak{h}^{i}(A)$, see [26, Theorem 5.1]. We are chiefly interested in the components $\mathfrak{h}^{3}(A)$ and $\mathfrak{h}^{2g-3}(A)$ when $g\geq 2$; in that case the Lefschetz decomposition has the form $\mathfrak{h}^{3}(A)=\mathfrak{h}^{3}_{\mathrm{prim}}(A)\oplus\ell\cdot\mathfrak{h}^{1}(A)$ and $\mathfrak{h}^{2g-3}(A)=\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)\oplus\ell^{g-2}\cdot\mathfrak{h}^{1}(A)$. It has the property that the isomorphism $\ell^{g-3}\colon\mathfrak{h}^{3}(A)\rightarrow\mathfrak{h}^{2g-3}(A)(g-3)$ is a direct sum of isomorphisms $\mathfrak{h}^{3}_{\mathrm{prim}}(A)\rightarrow\mathfrak{h}^{2g-3}_{\mathrm{prim}}(A)$ and $\ell\cdot\mathfrak{h}^{1}(A)\rightarrow\ell^{g-2}\mathfrak{h}^{1}(A)$. Taking Chow groups, we get a decomposition ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(A))={\rm CH}^{g-1}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(A))\oplus{\rm CH}^{g-1}(\ell^{g-2}\cdot\mathfrak{h}^{1}(A))$.
2.6. Beauville components of $C$
Consider a nice curve $C$ of genus $g\geq 2$ over $k$ with Jacobian $J$. Let $e$ be a degree-$1$ divisor on $C$ and embed $C$ in $J$ using the Abel–Jacobi map based at $e$, sending $x\in C$ to the divisor class of $x-e$. Decompose $[C]=[C]_{0}+\cdots+[C]_{g-1}$ with $[C]_{s}\in{\rm CH}_{(s)}^{g-1}(J)$. In this subsection we analyze the component $[C]_{1}$ more closely, whose vanishing is equivalent to the vanishing of $\kappa(C)$. Proposition 2.7 can be viewed as a determination of the ‘imprimitive part’ of $[C]_{1}$.
For $\alpha\in{\rm CH}_{p}(J)$ and $\beta\in{\rm CH}_{q}(J)$, the Pontryagin product $\alpha\star\beta$ is the pushforward of $\alpha\times\beta$ under the addition map $J\times J\rightarrow J$. If $n$ is a positive integer let $\alpha^{\star n}$ be the $n$-fold Pontryagin product of $\alpha$ with itself. Let $K_{C}\in{\rm CH}_{0}(C)$ denote the canonical divisor class and let $x_{e}\in J(k)$ be the point corresponding to the degree-00 divisor class $[(2g-2)e]-K_{C}$.
Proposition 2.7.
We have an equality in ${\rm CH}_{(1)}^{1}(J)$:
(2.3) $(2g-2)\cdot[C]_{0}^{\star(g-2)}\star[C]_{1}=[C]^{\star(g-1)}\star([0]-[x_{e}]).$
(Here we view $[0]-[x_{e}]$ as an element of ${\rm CH}_{0}(J)$.)
Proof.
If $D$ is a degree $g-1$ divisor class on $C$, let $\Theta_{D}$ be the image of the map ${\rm Sym}^{g-1}(C)\rightarrow J$ defined by $x\mapsto[x]-D$. Then $[C]^{\star(g-1)}=(g-1)![\Theta_{(g-1)e}]$. By Riemann–Roch,
$(-1)_{*}[\Theta_{(g-1)e}]=[\Theta_{K_{C}-(g-1)e}]=[x_{e}]\star[\Theta_{(g-1)e}].$
Combining the last two sentences shows that $(-1)_{*}([C]^{\star(g-1)})=[x_{e}]\star[C]^{\star(g-1)}$.
Since Pontryagin product sends ${\rm CH}^{g-p}_{(s)}(J)\times{\rm CH}^{g-q}_{(t)}(J)$ to ${\rm CH}^{g-p-q}_{(s+t)}(J)$ and since ${\rm CH}^{1}_{(s)}(J)\neq 0$ only if $s\in\{0,1\}$, we calculate that $[C]^{\star(g-1)}=[C]_{0}^{\star(g-1)}+(g-1)\cdot[C]_{0}^{\star(g-2)}\star[C]_{1}$. Applying $(-1)_{*}$ and $[x_{e}]\star$ to the previous identity, we obtain
$(-1)_{*}[C]^{\star(g-1)}$ $=[C]_{0}^{\star(g-1)}-(g-1)\cdot[C]_{0}^{\star(g-2)}\star[C]_{1},$
$[x_{e}]\star[C]^{\star(g-1)}$ $=[x_{e}]\star[C]_{0}^{\star(g-1)}+(g-1)\cdot[C]_{0}^{\star(g-2)}\star[C]_{1}.$
Note that $[x_{e}]\star$ acts trivially on ${\rm CH}^{1}_{(1)}(J)$ since $([x_{e}]-[0])\in\oplus_{s\geq 1}{\rm CH}^{g}_{(s)}(J)$ by the explicit description of the Beauville decomposition for zero-cycles [5, bottom of p. 649]. Equating the right hand sides of the two centered equations proves that $(2g-2)\cdot[C]_{0}^{\star(g-2)}\star[C]_{1}=[C]_{0}^{\star(g-1)}\star([0]-[x_{e}])$. Since $[C]_{s}^{\star(g-1)}\star([0]-[x_{e}])\in\oplus_{t\geq s+1}{\rm CH}^{1}_{(t)}(J)=\{0\}$ for all $s\geq 1$, it follows that $[C]_{0}^{\star(g-1)}\star([0]-[x_{e}])=[C]^{\star(g-1)}\star([0]-[x_{e}])$, concluding the proof. ∎
Corollary 2.8.
Suppose that $[C]_{1}=0$ in ${\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$. Then $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$.
Proof.
If $[C]_{1}=0$, then $[C]^{\star(g-1)}\star([0]-[x_{e}])=0$ by (2.3). On the other hand, $[C]^{\star(g-1)}=(g-1)![\Theta_{(g-1)e}]$ is multiple of a theta divisor, in the notation of the proof of Proposition 2.7. Since $\Theta_{(g-1)e}$ defines a principal polarization, the map $x\mapsto[\Theta_{(g-1)e+x}]-[\Theta_{(g-1)e}]=[\Theta_{(g-1)e}]\star([x]-[0])$ induces an isomorphism $\varphi\colon J(k)\rightarrow{\rm CH}^{1}(J)_{\hom}$. Since $\varphi(x_{e})$ is torsion, it follows that $x_{e}\in J(k)$ is itself torsion, as desired. ∎
The principal polarization defines an ample class $\ell\in{\rm CH}_{(1)}^{1}(J)$, which induces a Lefschetz decomposition $\mathfrak{h}^{2g-3}(J)=\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)\oplus\ell^{g-2}\cdot\mathfrak{h}^{1}(J)$ as in §2.5.
Corollary 2.9.
Suppose that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$. Then $[C]_{1}\in{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J))$.
Proof.
By definition and the discussion in §2.4, $[C]_{1}\in{\rm CH}_{(1)}^{g-1}(J)={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))$. Since $x_{e}$ is torsion, $(n)_{*}([0]-[x_{e}])=0$ for some integer $n\geq 1$. The decomposition (2.2) implies that $[0]=[x_{e}]$ in ${\rm CH}_{0}(J)_{{\rm\mathbb{Q}}}$. Therefore (2.3) shows that $[C]_{0}^{\star(g-2)}\star[C]_{1}=0$. Using properties of the $\mathfrak{sl}_{2}$-action on ${\rm CH}^{*}(J)_{{\rm\mathbb{Q}}}$ (in the sense of [33, §1.3]), this implies that $\ell\cdot[C]_{1}=0$. Since ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J))$ equals the kernel of $\ell\cdot(-)\colon{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J))\rightarrow{\rm CH}^{g-1}(\mathfrak{h}^{2g-1}(J)(1))$, the corollary follows. ∎
2.7. Ceresa cycles
Let $C/k$ be a nice curve of genus $g\geq 2$ with Jacobian $J$. Let $e$ be a degree-$1$ divisor on $C$ and let $\iota_{e}\colon C\rightarrow J$ be the Abel–Jacobi map based at $e$. We define $\kappa_{C,e}\in{\rm CH}_{1}(J)$ using the formula (1.1) from the introduction. Using the Beauville decomposition (2.2) to write $[C]=[\iota_{e}(C)]=\sum_{s=0}^{g-1}[C]_{s}$ with $[C]_{s}\in{\rm CH}^{g-1}_{(s)}(J)$, we calculate that
(2.4) $\kappa_{C,e}=\kappa(C)=2[C]_{1}+2[C]_{3}+\dots+2[C]_{2\lfloor\frac{g-2}{2}\rfloor+1}$
in ${\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$.
Lemma 2.10.
If $\kappa_{C,e}$ is torsion, then $(2g-2)e-K_{C}$ is torsion.
Proof.
If $\kappa_{C,e}$ is torsion, then $[C]_{1}=0$ by (2.4). We conclude using Corollary 2.8. ∎
Lemma 2.11.
If $e,e^{\prime}\in{\rm CH}_{0}(C)$ are degree-$1$ divisors such that $e-e^{\prime}$ is torsion, then $[\iota_{e}(C)]=[\iota_{e^{\prime}}(C)]$ and $\kappa_{C,e}=\kappa_{C,e^{\prime}}$ in ${\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$.
Proof.
Suppose $e-e^{\prime}$ has order $n$ in ${\rm CH}_{0}(C)$. Then $(n)\circ\iota_{e}=(n)\circ\iota_{e^{\prime}}$, hence $(n)_{*}([\iota_{e}(C)]-[\iota_{e^{\prime}}(C)])=0$ in ${\rm CH}_{1}(J)$. On the other hand, the decomposition (2.2) shows that $(n)_{*}\colon{\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$ is an isomorphism. Therefore $[\iota_{e}(C)]-[\iota_{e^{\prime}}(C)]$ is torsion. Hence $\kappa_{C,e}-\kappa_{C,e^{\prime}}$ is torsion too. ∎
Let $\kappa(C)$ be the image of $\kappa_{C,e}$ in ${\rm CH}_{1}(J)_{\rm\mathbb{Q}}$ for any choice of degree-$1$ divisor $e$ on $C$ such that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$. Lemma 2.11 shows that this class is independent of the choice of $e$. (If $k$ is not algebraically closed and no such $e$ exists over $k$, there exists a unique class $\kappa(C)\in{\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$ such that $\kappa(C)_{\bar{k}}=\kappa(C_{\bar{k}})$ in ${\rm CH}_{1}(J_{\bar{k}})_{{\rm\mathbb{Q}}}$, since Chow groups with ${\rm\mathbb{Q}}$-coefficients satisfy Galois descent.) We let $\bar{\kappa}(C)$ be the image of $\kappa(C)$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$. Since all degree-$1$ divisors $e$ on $C$ are algebraically equivalent, $\bar{\kappa}(C)$ is also the image of $\kappa_{e}(C)$ in ${\rm Gr}_{1}(J)_{\rm\mathbb{Q}}$ for any degree-$1$ divisor, not necessarily with the property that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$.
Suppose now that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$. Since $[\iota_{e}(C)]\in{\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$ is independent of the choice of $e$, the same is true for the classes $[C]_{s}$. In particular, they are ${\rm Aut}(C)$-invariant.
Proposition 2.12.
In the above notation, $\kappa(C)=0$ in ${\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$ if and only if $[C]_{1}=0$ in ${\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$ if and only if $[C]_{s}=0$ for all $s\geq 1$. Moreover $\bar{\kappa}(C)=0$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ if and only if $[C]_{1}\in{\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ if and only if $[C]_{s}=0$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ for all $s\geq 1$.
Proof.
Using the expression (2.4), it suffices to prove $[C]_{1}=0$ implies $[C]_{s}=0$ for all $s\geq 1$. This follows from the third centered equation of [46, Theorem 1.5.5]. The proof for $\bar{\kappa}(C)$ is identical. ∎
Proposition 2.12 shows that the vanishing of $\kappa(C)$ is equivalent to the vanishing of $[C]_{1}$. The next lemma isolates a summand of the Chow group that contains $[C]_{1}$. Recall that if $M$ is a direct summand of $\mathfrak{h}(J)$, then ${\rm CH}^{p}(M)$ is naturally a summand of ${\rm CH}^{p}(\mathfrak{h}(J))={\rm CH}^{p}(J)_{{\rm\mathbb{Q}}}$.
Proposition 2.13.
In the above notation, $[C]_{1}\in{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)^{{\rm Aut}(C)})\subset{\rm CH}_{1}(J)_{{\rm\mathbb{Q}}}$.
Proof.
Combine Corollary 2.9 and the fact that $[C]_{1}$ is ${\rm Aut}(C)$-invariant. ∎
3. Vanishing criteria in the Chow and Griffiths groups
The proof of Theorem A is quite short, but uses in a crucial way Kimura’s notion of finite dimensional Chow motives [25].
Proof of Theorem A.
We use the definitions and notations of §2.3 and §2.7. Choose a degree-$1$ divisor $e$ such that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)_{{\rm\mathbb{Q}}}$ and decompose $[\iota_{e}(C)]=\sum_{s=-1}^{g-1}[C]_{s}$ with $[C]_{s}\in{\rm CH}^{g-1}_{(s)}(J)$. Let $G={\rm Aut}(C)$. Proposition 2.13 shows that $[C]_{1}\in{\rm CH}^{g-1}(\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)^{G})$.
By the hypothesis and the Lefschetz isomorphism, $\operatorname{H}^{*}(\mathfrak{h}_{\mathrm{prim}}^{2g-3}(J)^{G})=\operatorname{H}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$. Since any summand of the motive of an abelian variety is finite-dimensional in the sense of Kimura [25, Example 9.1], it follows from Kimura’s [25, Corollary 7.3] that $\mathfrak{h}^{2g-3}_{\mathrm{prim}}(J)^{G}=0$ and hence $[C]_{1}=0$ in ${\rm CH}^{g-1}(J)_{{\rm\mathbb{Q}}}$. By Proposition 2.12, we conclude that $\kappa(C)=0$. ∎
Example 3.1.
There is exactly one non-hyperelliptic genus 3 curve over ${\rm\mathbb{C}}$ for which the criterion of [39] applies, namely the curve $y^{3}=x^{4}+1$. The curve $y^{3}=x^{4}+x$ satisfies the weaker hypothesis of Theorem A (as was observed in [7]), so we deduce that $\kappa(C)=0$ for this curve as well. Beauville and Schoen studied the specific geometry of this curve and showed that $\bar{\kappa}(C)=0$ [8].
Example 3.2.
The genus $4$ curve $y^{3}=x^{5}+1$ satisfies $\operatorname{H}^{3}(J)^{{\rm Aut}(C)}=0$ [29, proof of Theorem 3.3], so $\kappa(C)=0$.
Theorem B will follow from the next proposition. In that proposition and its proof, if $X$ is a nice variety over ${\rm\mathbb{C}}$ we write $\operatorname{H}^{*}(X)$ for the singular cohomology of $X({\rm\mathbb{C}})$ with ${\rm\mathbb{Q}}$-coefficients, seen as an object in the category of Hodge structures.
Proposition 3.3.
Let $C$ be a smooth, projective, integral curve over ${\rm\mathbb{C}}$ with Jacobian $J$ and let $G\subset{\rm Aut}(C)$ be a subgroup with $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{G}=0$. Then there exists an abelian variety $A/{\rm\mathbb{C}}$ such that $\operatorname{H}^{3}(J)^{G}\simeq\operatorname{H}^{1}(A)(-1)$. If the Hodge conjecture holds for $J\times A$, then $\mathfrak{h}^{3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-1)$ and $\bar{\kappa}(C)=0$.
Proof.
Let $N^{1}\operatorname{H}^{3}(J)$ be the largest sub-Hodge structure of $\operatorname{H}^{3}(J)$ of type $(1,2)+(2,1)$. The polarization on $\operatorname{H}^{1}(J)$ induces a polarization on $N^{1}\operatorname{H}^{3}(J)$ and so $N^{1}\operatorname{H}^{3}(J)\simeq\operatorname{H}^{1}(B)(-1)$ for some abelian variety $B/{\rm\mathbb{C}}$. The assumptions and the Hodge decomposition imply that $\operatorname{H}^{3}(J)^{G}$ is a Hodge structure of type $(1,2)+(2,1)$. It follows that $\operatorname{H}^{3}(J)^{G}\subset N^{1}\operatorname{H}^{3}(J)$, so there exists an abelian subvariety $A\subset B$ with $\operatorname{H}^{3}(J)^{G}\simeq\operatorname{H}^{1}(A)(-1)$.
We now show that the Hodge conjecture for $J\times A$ implies the claims of the final sentence. Fix mutually inverse isomorphisms of Hodge structures $\phi\colon\operatorname{H}^{3}(J)^{G}\rightarrow\operatorname{H}^{1}(A)(-1)$ and $\psi\colon\operatorname{H}^{1}(A)(-1)\rightarrow\operatorname{H}^{3}(J)$. By the Hodge conjecture, there exist morphisms of motives $\Phi\colon\mathfrak{h}^{3}(J)^{G}\rightarrow\mathfrak{h}^{1}(A)(-1)$ and $\Psi\colon\mathfrak{h}^{1}(A)(-1)\rightarrow\mathfrak{h}^{3}(J)^{G}$ (in other words, cycles on $J\times A$ with certain properties) with $\operatorname{H}^{*}(\Phi)=\phi$ and $\operatorname{H}^{*}(\Psi)=\psi$. Since $\mathfrak{h}^{1}(A)(-1)$ and $\mathfrak{h}^{3}(J)^{G}$ are Kimura finite-dimensional [25, Example 9.1] and $\operatorname{H}^{*}(\Phi)\circ\operatorname{H}^{*}(\Psi)$ and $\operatorname{H}^{*}(\Psi)\circ\operatorname{H}^{*}(\Phi)$ are the identity, it follows from [25, Proposition 7.2(ii)] (see also [4, Corollaire 3.16]) that $\Psi\circ\Phi$ and $\Phi\circ\Psi$ are themselves isomorphisms. Hence $\Phi$ and $\Psi$ are isomorphisms too and we conclude that $\mathfrak{h}^{3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-1)$.
Using the Lefschetz isomorphism $\mathfrak{h}^{2g-3}(J)\simeq\mathfrak{h}^{3}(J)(-g+3)$ of [26, Theorem 5.2], we obtain an isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$. Similarly to the proof of Theorem A, we decompose $[\iota_{e}(C)]=\sum_{s=0}^{g-1}[C]_{s}$ with $[C]_{s}\in{\rm CH}^{g-1}_{(s)}(J)$ and observe that the class $[C]_{1}$ lies in ${\rm CH}_{(1)}^{g-1}(J)^{G}={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})$. Lemma 2.6 combined with the isomorphism $\mathfrak{h}^{2g-3}(J)^{G}\simeq\mathfrak{h}^{1}(A)(-g+2)$ shows that ${\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})={\rm CH}^{g-1}(\mathfrak{h}^{2g-3}(J)^{G})_{\mathrm{alg}}$, hence every element of ${\rm CH}^{g-1}_{(1)}(J)^{G}$ lies in ${\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$. Therefore $[C]_{1}\in{\rm CH}^{g-1}(J)_{\mathrm{alg},{\rm\mathbb{Q}}}$, hence the image of $[C]_{1}$ in ${\rm Gr}_{1}(J)_{{\rm\mathbb{Q}}}$ vanishes. Proposition 2.12 then implies that $\bar{\kappa}(C)$ vanishes too. ∎
Proof of Theorem B.
Since $C$ can be defined over a countable field and since such a field can be embedded in ${\rm\mathbb{C}}$, Lemma 2.1 shows that we may assume $k={\rm\mathbb{C}}$. We conclude by Proposition 3.3 applied to $G={\rm Aut}(C)$. ∎
Example 3.4.
Let $C\colon y^{3}=x^{4}+ax^{2}+bx+c$ be a Picard curve. Then $\operatorname{H}^{0}(C,\Omega_{C}^{1})\simeq\chi\oplus\chi\oplus\chi^{2}$ as $C_{3}$-representations, where $\chi$ is a character of order $3$. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{{\rm Aut}(C)}=0$, so that the condition of Theorem B is satisfied. We consider these curves in detail in the next section.
We exhibit two families of examples of Theorem B in genus $4$. It would be interesting to study these families in more detail (along the lines of what we do in the next section for Picard curves).
Example 3.5.
Consider for every $t\in{\rm\mathbb{C}}\setminus\{0,1,-1\}$ the $\mu_{5}$-cover of ${\rm\mathbb{P}}^{1}$ with equation
$C\colon y^{5}=x^{3}(x-1)^{2}(x-t).$
By [34, Lemma 2.7], the $\mu_{5}$-representation $\operatorname{H}^{0}(C,\Omega_{C})$ is the direct sum of the four nontrivial characters. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{\mu_{5}}=0$ and $\operatorname{H}^{3}(J)^{\mu_{5}}$ is isomorphic to $\operatorname{H}^{1}(A)(-1)$, for some abelian surface $A$ (using Proposition 3.3). We have $D_{5}\subset{\rm Aut}(C)$ by [22, Corollary 1]. By the Kani-Rosen formula [23, Theorem B], the Jacobian $J$ is isogenous to a product of abelian surfaces of the form $\operatorname{Jac}(C/\tau)$, for some involution $\tau$. The Hodge conjecture in any codimension is known for products of abelian surfaces [40, Theorem 3.15], so $\bar{\kappa}(C)=0$ by Proposition 3.3.
Example 3.6.
The general $\mu_{3}$-cover of ${\rm\mathbb{P}}^{1}$ with equation
$y^{3}=x^{2}(x-1)^{2}(x^{3}+ax^{2}+bx+c),$
has genus $4$ and there is an isomorphism of $\mu_{3}$-representations $\operatorname{H}^{0}(C,\Omega_{C})\simeq\chi\oplus\chi\oplus\chi^{2}\oplus\chi^{2}$. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{\mu_{3}}=0$, and the abelian variety $A$ from Proposition 3.3 is 4-dimensional. If the Hodge conjecture holds for the $8$-dimensional $J\times A$, then $\bar{\kappa}(C)=0$.
4. Picard curves
4.1. Generalities
Let $k$ be a field of characteristic zero. A Picard curve over $k$ is by definition a nice curve with an affine model $y^{3}=f(x)=x^{4}+ax^{2}+bx+c$ for some $a,b,c\in k$. Conversely, given such a polynomial $f(x)\in k[x]$ of nonzero discriminant, the projective closure of $y^{3}=f(x)$ in ${\rm\mathbb{P}}^{2}_{k}$ is a nice curve denoted by $C_{f}$. It has a unique point at infinity $P_{\infty}$, which is $k$-rational.
For every third root of unity $\omega\in\bar{k}$, the map $(x,y)\mapsto(x,\omega y)$ defines an automorphism of $C_{f,\bar{k}}$. We view $\mu_{3}$ as a subgroup of ${\rm Aut}(C_{f,\bar{k}})$ in this way. Then $\mu_{3}$ also acts on the Jacobian $J_{f,\bar{k}}$ by taking images of divisors.
The discriminant of $f$ has the following expression:
(4.1) ${\rm disc}(f)=-4a^{3}b^{2}-27b^{4}+16a^{4}c+144ab^{2}c-128a^{2}c^{2}+256c^{3}.$
We view $f$ as the dehomogenization $F(x,1)$ of the quartic form $F(X,Z)=X^{4}+aX^{2}Z^{2}+bXZ^{3}+cZ^{4}$, and we define $I(f)$ and $J(f)$ to be the usual $I$- and $J$-invariants attached to $F$, as in [10, §2]. Their explicit formulae in our case are:
$I(f)$ $=a^{2}+12c,$
$J(f)$ $=72ac-2a^{3}-27b^{2}.$
The 19th century invariant theorists observed the identity $J(f)^{2}=4I(f)^{3}-27\cdot{\rm disc}(f)$, which can be verified by direct computation. Therefore $P_{f}:=(I(f),J(f))$ is a $k$-point on the elliptic curve
$E_{f}\colon y^{2}=4x^{3}-27\cdot{\rm disc}(f).$
4.2. Ceresa vanishing criteria
Since $(2g-2)P_{\infty}=4P_{\infty}$ is canonical, we may use $P_{\infty}$ to embed $C_{f}$ in its Jacobian $J_{f}$ and define the Ceresa cycle $\kappa_{C_{f},P_{\infty}}\in{\rm CH}_{1}(J_{f})$ as in the introduction; we denote it by $\kappa_{f}$ for simplicity. Recall that $\kappa(C_{f})$ denotes the image of $\kappa_{f}$ in ${\rm CH}_{1}(J_{f})_{{\rm\mathbb{Q}}}$ and $\bar{\kappa}(C_{f})$ its image in ${\rm Gr}_{1}(J_{f})_{{\rm\mathbb{Q}}}$. Theorem C follows from the following slightly stronger theorems, whose proofs will take up the rest of this section.
Theorem 4.1.
There exists an integer $N\geq 1$ ((((depending neither on $f$ nor $k)$ such that $N\cdot\kappa_{f}\in{\rm CH}^{2}(J_{f})_{\mathrm{alg}}$ for every Picard curve $C_{f}$ over every algebraically closed field $k$ (of characteristic zero).
Theorem 4.2.
The Ceresa cycle $\kappa_{f}\in{\rm CH}_{1}(J)$ is torsion if and only if $P_{f}\in E_{f}(k)$ is torsion. Moreover, if $k$ is algebraically closed, there exists an integer $M\geq 1$ with the following property: if $C_{f}$ is a Picard curve and $\kappa_{f}$ is torsion, then $\operatorname*{ord}(\kappa_{f})$ divides $M\cdot\operatorname*{ord}(P_{f})$ and $\operatorname*{ord}(P_{f})$ divides $M\cdot\operatorname*{ord}(\kappa_{f})$.
Theorem 4.1 will be proven in §4.4, and Theorem 4.2 will be proven in §4.7. A standard argument using Lemma 2.1 shows that we may assume $k={\rm\mathbb{C}}$. So in the remainder of §4, all varieties will be over ${\rm\mathbb{C}}$, and cohomology will be singular cohomology.
Remark 4.3.
Theorem 4.2 generalizes [27, Theorem 5.16], which considered the special case where $b=0$. There, we exploited the bielliptic cover to show that the Ceresa cycle maps via a correspondence to a multiple of the point $Q_{f}=(a^{2}-4c,a(a^{2}-4c))$ on the elliptic curve $E_{f}^{\prime}\colon y^{2}=x^{3}+16\cdot{\rm disc}(f)$. This is compatible with the general case since there is a $3$-isogeny $\phi_{f}\colon E^{\prime}_{f}\to E_{f}$, and one checks using the explicit formula for $\phi_{f}$ [9, Equation (2)] that $\phi_{f}(Q_{f})=P_{f}$.
Remark 4.4.
Is it always the case that $\kappa_{f}\neq 0$? (Recall that $\kappa_{f}$ lies in the Chow group with ${\rm\mathbb{Z}}$-coefficients.) We cannot conclude this from our proof of Theorem 4.2 below since we have worked with ${\rm\mathbb{Q}}$-coefficients, and we make use of various isogenies whose degrees we do not control.
4.3. Multilinear algebra
Our first goal (Proposition 4.7) is to explicitly identify the abelian variety $A$ of Proposition 3.3 for Picard curves.
Write $\mathcal{O}={\rm\mathbb{Z}}[\omega]$ for the ring of Eisenstein integers with $\omega^{2}+\omega+1=0$ and let $K={\rm\mathbb{Q}}(\sqrt{-3})$ be its fraction field. Let $C$ be a Picard curve over ${\rm\mathbb{C}}$ with Jacobian variety $J$. The $\mu_{3}$-action on $C$ extends to an embedding $\mathcal{O}\subset{\rm End}(J)$. Using this action, the singular cohomology group $\operatorname{H}^{1}(J;{\rm\mathbb{Z}})$ is a free $\mathcal{O}$-module of rank $3$, and $\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ is a $3$-dimensional $K$-vector space. The next lemma says that the criterion of Theorem B is always satisfied for Picard curves.
Lemma 4.5.
$\operatorname{H}^{0}(J,\Omega_{J}^{3})^{\mu_{3}}=0$ and the Hodge structure $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}$ is of type $(1,2)+(2,1)$.
Proof.
Since $\operatorname{H}^{0}(J,\Omega_{J}^{3})\simeq\bigwedge^{3}\operatorname{H}^{0}(C,\Omega_{C}^{1})$, the first claim follows from a calculation with differentials (Example 3.4). The second claim follows from the Hodge decomposition for $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$. ∎
We may view $\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ either as a $K$-vector space or ${\rm\mathbb{Q}}$-vector space; when we perform tensor operations, we will add the subscript $K$ when we view it as a $K$-vector space, and add no subscript otherwise. For example, cup product induces an isomorphism $\bigwedge^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})\simeq\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$, and we will use this identification without further mention.
The universal property of exterior powers induces a canonical ${\rm\mathbb{Q}}$-linear surjection $\bigwedge^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})\rightarrow\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$. It is well known (see [35, Lemma 12(i)] or [17, Lemma 4.3]) that this map admits a canonical splitting, which we use to view $\bigwedge_{K}^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ as a direct summand of $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$.
Lemma 4.6.
We have $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}=\bigwedge_{K}^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ inside $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$. Moreover $\dim_{{\rm\mathbb{Q}}}\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}=2$.
Proof.
It suffices to prove the statements after tensoring with ${\rm\mathbb{C}}$. Let $g\in\mu_{3}\subset{\rm Aut}(C)$ be a nontrivial element. The action of $g$ on $\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ has eigenvalues (with multiplicity) $\omega,\omega,\omega,\omega^{2},\omega^{2},\omega^{2}$ so we can write $\operatorname{H}^{1}(J;{\rm\mathbb{C}})=V_{1}\oplus V_{2}$ where $V_{i}$ is the $\omega^{i}$-eigenspace. Since a three element subset of these eigenvalues have product $1$ if and only if they are all equal, $\operatorname{H}^{3}(J;{\rm\mathbb{C}})^{\mu_{3}}=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$. On the other hand, an argument similar to the proof of [17, Proposition 4.4] shows $(\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}}))\otimes_{K}{\rm\mathbb{C}}=\bigwedge^{3}_{K\otimes{\rm\mathbb{C}}}\operatorname{H}^{1}(J;{\rm\mathbb{C}})=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$, proving the equality. The explicit description of this subspace shows that it is $2$-dimensional over ${\rm\mathbb{Q}}$. ∎
Let $E$ be the elliptic curve with Weierstrass equation $y^{2}=x^{3}+1$.
Proposition 4.7.
There is an isomorphism of Hodge structures $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}\simeq\operatorname{H}^{1}(E,{\rm\mathbb{Q}})(-1)$.
Proof.
Since there exists a unique elliptic curve up to isogeny with endomorphism algebra $K$, there exists a unique Hodge structure of dimension $2$, type $(0,1)+(1,0)$ and carrying an action of $K$. Since both $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}$ and $\operatorname{H}^{1}(E;{\rm\mathbb{Q}})$ have these properties (the former by Lemma 4.6), they must be isomorphic, and we conclude using Lemma 4.6. ∎
4.4. Weil classes and Schoen’s theorem
Our next goal is to upgrade the isomorphism of Proposition 4.7 to an isomorphism in the category of Chow motives, and hence deduce (using Proposition 3.3) the vanishing of $\bar{\kappa}(C)$ in the Griffiths group. We use the following special case of a result of Schoen, which crucially uses the assumption that the endomorphism algebra of $J\times E$ contains $K={\rm\mathbb{Q}}(\omega)$:
Theorem 4.8 (Schoen).
The Hodge conjecture holds for $J\times E$.
Proof.
Since $J\times E$ is four-dimensional, it suffices to prove Hodge classes in $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. By [36, Theorem (0.1), Part (i) and (iv)], such Hodge classes are sums of products of divisor classes (which are algebraic) and Weil classes $W_{K}:=\bigwedge^{4}_{K}\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})\subset\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$. Since $K={\rm\mathbb{Q}}(\omega)$ and the embedding $K\subset{\rm End}(J\times E)\otimes{\rm\mathbb{Q}}$ can be chosen to have signature $(2,2)$, Schoen has shown in [41] that the classes in $W_{K}$ are algebraic, concluding that all Hodge classes of $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. ∎
Recall from §2.3 our conventions on motives, the canonical Chow–Künneth components $\mathfrak{h}^{i}(J)$ and $\mathfrak{h}^{j}(E)$, and the motive of fixed points of a finite group action.
Corollary 4.9.
There is an isomorphism $\mathfrak{h}^{3}(J)^{\mu_{3}}\simeq\mathfrak{h}^{1}(E)(-1)$.
Proof.
This follows from Proposition 3.3, using Proposition 4.7 and Theorem 4.8. ∎
Corollary 4.10.
If $C$ is a Picard curve over ${\rm\mathbb{C}}$, then $\bar{\kappa}(C)=0$.
Proof.
This follows from Proposition 3.3, using Proposition 4.7 and Theorem 4.8. ∎
Proof of Theorem 4.1.
Consider the parameter space $S_{0}=\{(a,b,c)\mid{\rm disc}(f)\neq 0\}\subset{\rm\mathbb{A}}^{3}_{{\rm\mathbb{Q}}}$ of Picard curves. Let $\eta$ be the generic point of $S_{0}$ with function field $k(\eta)={\rm\mathbb{Q}}(a,b,c)$, let $C_{\eta}$ be the generic Picard curve over $k(\eta)$ with Jacobian $J_{\eta}$, and let $\kappa_{C_{\eta}}\in{\rm CH}_{1}(J_{\eta})$ be the Ceresa cycle of $C_{\eta}$ based at the point at infinity. Fix an embedding $j\colon k(\eta)\rightarrow{\rm\mathbb{C}}$. By Corollary 4.10, the base change of $\kappa_{C{\eta}}$ along $j$ is torsion in the Griffiths group. By Lemma 2.1, this implies $\kappa_{C_{\eta}}$ is itself torsion in ${\rm Gr}_{1}(J_{\eta})$. Let $N_{0}$ be the finite order of $\kappa_{C_{\eta}}$ in ${\rm Gr}_{1}(J_{\eta})$. By spreading out, it follows that $\kappa_{C_{f}}$ is $N_{0}$-torsion in ${\rm Gr}_{1}(J_{f})$ for all $f$ in an open dense subset $U_{0}\subset S_{0}$. By applying the same argument to the generic points of the irreducible components of $S_{0}\setminus U_{0}$, there exists an open dense $U_{1}\subset S_{0}$ containing $U_{0}$ and those generic points such that $\kappa_{C_{f}}$ is $N_{1}$-torsion in ${\rm Gr}_{1}(J_{f})$ for all $f$ in $U_{1}$. Repeating this process, we obtain a sequence of open subsets $U_{0}\subset U_{1}\subset\cdots\subset S_{0}$ whose complements have strictly decreasing codimension. Therefore this sequence must terminate after finitely many steps hence there exists an integer $N$ such that $\kappa_{C_{f}}$ is $N$-torsion in ${\rm Gr}_{1}(J_{f})$ for all $f\in S_{0}(k)$ over every algebraically closed field of characteristic zero. ∎
The remainder of the section is devoted to proving Theorem 4.2. To this end, we will analyze the Abel–Jacobi image of $\kappa_{f}$ (the “normal function” associated to $\kappa_{f}$) in the next two subsections.
4.5. The Abel–Jacobi map
If $X$ is a smooth variety over ${\rm\mathbb{C}}$, we will use the notion of an integral (respectively rational) variation of (pure) Hodge structures over the complex manifold $X({\rm\mathbb{C}})$, called a ${\rm\mathbb{Z}}$-VHS (respectively ${\rm\mathbb{Q}}$-VHS) for short; see [45, §5.3.1] for definitions. If $V$ is a Hodge structure of odd weight $12k-$ 1, its intermediate Jacobian $\mathrm{J}(V)$ is the complex torus
$(V\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{C}})/(F^{k}+V_{\tau}),$
where $F^{k}\subset V\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{C}}$ is a part of the descending Hodge filtration and $V_{\tau}$ denotes the quotient of $V$ by its torsion subgroup. More generally, if $\mathbb{V}$ is a ${\rm\mathbb{Z}}$-VHS of weight $12k-$ 1 over $X({\rm\mathbb{C}})$, we can define its intermediate Jacobian $\mathrm{J}(\mathbb{V})\rightarrow X({\rm\mathbb{C}})$, a relative complex torus whose fibers over points $x\in X({\rm\mathbb{C}})$ are the classical intermediate Jacobians $\mathrm{J}(\mathbb{V}_{x})$, see [45, §7.1.1]. If $\mathbb{V},\mathbb{W}$ are two ${\rm\mathbb{Z}}$-VHS of odd weight, then $\mathrm{J}(\mathbb{V}(p))=\mathrm{J}(\mathbb{V})$ for all $p\in{\rm\mathbb{Z}}$ and a morphism $\mathbb{V}\rightarrow\mathbb{W}$ of ${\rm\mathbb{Z}}$-VHS induces a homomorphism of (relative) complex tori $\mathrm{J}(\mathbb{V})\rightarrow\mathrm{J}(\mathbb{W})$.
If $X/{\rm\mathbb{C}}$ is a nice variety and $0\leq p\leq\dim(X)$, we write $\mathrm{J}^{p}(X)=\mathrm{J}(\operatorname{H}^{2p-1}(X({\rm\mathbb{C}});{\rm\mathbb{Z}}))$. In this situation there is an Abel–Jacobi map
(4.2) $\mathrm{AJ}^{p}_{X}\colon{\rm CH}^{p}(X)_{\hom}\rightarrow\mathrm{J}^{p}(X),$
defined in [45, §7.2.1]. Moreover, if $S/{\rm\mathbb{C}}$ is a smooth variety, $\pi\colon X\rightarrow S$ a smooth projective morphism with geometrically integral fibres and $p\in{\rm\mathbb{Z}}_{\geq 0}$, then $R^{p}\pi_{*}{\rm\mathbb{Z}}$ (pushforward of the constant sheaf in the analytic topology) has the structure of a ${\rm\mathbb{Z}}$-VHS over $S({\rm\mathbb{C}})$. If $Z$ is a codimension $p$ cycle on $X$ all of whose components are flat over $S$, and such that $Z_{s}\in{\rm CH}^{p}(X_{s})_{\hom}$ for every $s\in S({\rm\mathbb{C}})$, then Griffiths has shown that there exists a holomorphic section ${\rm AJ}(Z)$ of the relative complex torus $\mathrm{J}(R^{p}\pi_{*}{\rm\mathbb{Z}})\rightarrow S({\rm\mathbb{C}})$ with the property that ${\rm AJ}(Z)_{s}={\rm AJ}_{X}^{p}(Z_{s})$ for all $s\in S({\rm\mathbb{C}})$; this is called the normal function associated to $Z$.
We record the fact that Abel–Jacobi maps are compatible with correspondences. Let $X,Y$ be nice varieties over ${\rm\mathbb{C}}$ and let $\gamma\in{\rm CH}^{r+\dim(X)}(X\times Y)$ be a correspondence of degree $r$. This induces for every $p\geq 0$ a homomorphism $\gamma_{*}\colon{\rm CH}^{p}(X)\rightarrow{\rm CH}^{p+r}(Y)$ via the formula $\alpha\mapsto\pi_{Y,*}(\pi_{X}^{*}(\alpha)\cdot\gamma)$, where $\pi_{X}\colon X\times Y\rightarrow X$ and $\pi_{Y}\colon X\times Y\rightarrow Y$ denote the projections. The same formula defines morphism of Hodge structures $\operatorname{H}^{p}(X;{\rm\mathbb{Z}})\rightarrow\operatorname{H}^{p+2r}(Y)(r)$ for every $p$, hence a homomorphism of complex tori $\gamma_{*}\colon\mathrm{J}^{p}(X)\rightarrow\mathrm{J}^{p+r}(Y)$.
Lemma 4.11.
In the above notation, $\gamma_{*}$ sends ${\rm CH}^{p}(X)_{\hom}$ to ${\rm CH}^{p+r}(Y)_{\hom}$. Moreover for every $\alpha\in{\rm CH}^{p}(X)_{\hom}$, $\gamma_{*}({\rm AJ}_{X}^{p}(\alpha))={\rm AJ}_{Y}^{p+r}(\gamma_{*}(\alpha))$.
Proof.
The first sentence follows from the compatibility of the cycle class map with correspondences [45, Proposition 9.21]. To prove the compatibility of the Abel–Jacobi map with correspondences, it suffices to prove the compatibility with pullbacks, pushforwards and intersection product. The case of pullback is elementary, using the definition of ${\rm AJ}^{p}_{X}$ in terms of extensions of Hodge structures, see [15, §2.2]. The case of pushforward follows from that of pullback and Poincare duality. Finally, compatibility with intersection product follows from [45, Proposition 9.23]. ∎
Let $S=\{(a,b,c)\mid{\rm disc}(f)\neq 0\}\subset{\rm\mathbb{A}}^{3}_{{\rm\mathbb{C}}}$ be the parameter space of Picard curves over ${\rm\mathbb{C}}$. We will identify ${\rm\mathbb{C}}$-valued points of $S$ with polynomials $f=x^{4}+ax^{2}+bx+c\in{\rm\mathbb{C}}[x]$ of nonzero discriminant. Let $\mathcal{C}\rightarrow S$ be the universal Picard curve, and let $\pi\colon\mathcal{J}\rightarrow S$ its relative Jacobian variety. The point at infinity defines a section $P_{\infty}$ of $\mathcal{C}$. Let $\kappa_{\mathcal{C}}\in{\rm CH}_{1}(\mathcal{J})$ be the universal Ceresa cycle with respect to this section. (Comparing with our earlier notation, we have $\mathcal{C}_{f}=C_{f}$ and $\kappa_{\mathcal{C},f}=\kappa_{f}$ for every $f\in S({\rm\mathbb{C}})$.)
Let $\mathbb{V}=R^{3}\pi_{*}{\rm\mathbb{Z}}$ be the ${\rm\mathbb{Z}}$-VHS on $S({\rm\mathbb{C}})$ interpolating the cohomology groups $\operatorname{H}^{3}(\mathcal{J}_{f};{\rm\mathbb{Z}})$ for $f\in S({\rm\mathbb{C}})$. Then the normal function ${\rm AJ}(\kappa_{\mathcal{C}})$ is a section of $\mathrm{J}(\mathbb{V})\rightarrow S({\rm\mathbb{C}})$ interpolating ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$.
Proposition 4.12.
For $f\in S({\rm\mathbb{C}})$, the class $\kappa_{f}\in{\rm CH}^{2}(J_{f})_{\mathrm{hom}}$ is torsion if and only if ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})\in\mathrm{J}^{2}(J_{f})$ is torsion. The torsion order of $\kappa_{f}$, if finite, equals the torsion order of ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$.
Proof.
Let $E$ be the elliptic curve $y^{2}=x^{3}+1$. Corollary 4.9 shows that there exists a correspondence $\gamma\in{\rm CH}^{2}(E\times J)$ such that $\gamma_{*}$ induces isomorphisms ${\rm CH}^{1}(E)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{2}(J)^{\mu_{3}}_{{\rm\mathbb{Q}}}$ and $\operatorname{H}^{1}(E;{\rm\mathbb{Q}})\rightarrow\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}(1)$. By Lemma 4.11 these form a commutative diagram:
[…]
All arrows except the horizontal bottom one are isomorphisms of abelian groups. The same is therefore true for the bottom one ${\rm AJ}\colon{\rm CH}^{2}(J_{f})^{\mu_{3}}_{\hom,{\rm\mathbb{Q}}}\rightarrow\mathrm{J}^{2}(J_{f})\otimes{\rm\mathbb{Q}}$. Since $\kappa_{f}\in{\rm CH}^{2}(J_{f})_{\hom}^{\mu_{3}}$, we conclude that $\kappa_{f}$ is torsion if and only if ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$ is. The last claim follows from the fact that ${\rm AJ}^{2}_{J_{f}}$ is injective on torsion subgroups by a result of Murre [37, Theorem 10.3]. ∎
4.6. Identifying the complex torus $\mathrm{J}(\mathbb{V}^{\mu_{3}})$
The $\mu_{3}$-action on $\mathcal{C}$ induces, via functoriality, a $\mu_{3}$-action on $\mathcal{J}$, $\mathbb{V}$ and $\mathrm{J}(\mathbb{V})$. The subsheaf of fixed points $\mathbb{V}^{\mu_{3}}$ has the structure of a ${\rm\mathbb{Z}}$-VHS. The connected component of the identity $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$ of $\mathrm{J}(\mathbb{V})^{\mu_{3}}$ is a relative complex torus over $S({\rm\mathbb{C}})$. Moreover, the natural homomorphism $\mathrm{J}(\mathbb{V}^{\mu_{3}})\rightarrow\mathrm{J}(\mathbb{V})$ induces an isomorphism onto $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$.
Lemma 4.13.
The multiple $3\cdot\mathrm{AJ}(\kappa_{\mathcal{C}})$ lands in $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$.
Proof.
Since the Ceresa cycle $\kappa_{\mathcal{C}}$ is $\mu_{3}$-invariant, $\mathrm{AJ}(\kappa_{\mathcal{C}})$ lands in $\mathrm{J}(\mathbb{V})^{\mu_{3}}$. The norm map $\mathrm{N}\colon\mathrm{J}(\mathbb{V})\rightarrow\mathrm{J}(\mathbb{V})$, defined by $x\mapsto x+\omega\cdot x+\omega^{2}\cdot x$, lands in $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$, since the image must be connected and $\mu_{3}$-invariant. We conclude that $\mathrm{N}(\mathrm{AJ}(\kappa_{\mathcal{C}}))=3\cdot\mathrm{AJ}(\kappa_{\mathcal{C}})$ lands in $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$. ∎
Therefore $3\cdot\mathrm{AJ}(\kappa_{\mathcal{C}})$ defines a section of $\mathrm{J}(\mathbb{V})^{\mu_{3},\circ}$, hence we may view it as a section of the complex torus $\mathrm{J}(\mathbb{V}^{\mu_{3}})\rightarrow S({\rm\mathbb{C}})$ in what follows. We study this relative complex torus (up to isogeny) in the next two propositions.
Proposition 4.14.
Let $\mathcal{E}\rightarrow S$ be the relative elliptic curve with Weierstrass equation
$\mathcal{E}\colon y^{2}=4x^{3}-27\cdot{\rm disc}(f).$
Then the relative complex torus $\mathrm{J}(\mathbb{V}^{\mu_{3}})\rightarrow S$ is isogenous to the relative complex torus $\mathcal{E}({\rm\mathbb{C}})\rightarrow S({\rm\mathbb{C}})$.
Proof.
Recall from §4.3 that we write $\mathcal{O}={\rm\mathbb{Z}}[\omega]$ and $K={\rm\mathbb{Q}}(\omega)$. The ${\rm\mathbb{Z}}$-VHS $R^{1}\pi_{*}{\rm\mathbb{Z}}$ interpolating the cohomology groups $\operatorname{H}^{1}(\mathcal{J}_{f};{\rm\mathbb{Z}})$ comes equipped with an action of $\mathcal{O}$, and cup product induces an isomorphism $\bigwedge^{3}R^{1}\pi_{*}{\rm\mathbb{Z}}\simeq R^{3}\pi_{*}{\rm\mathbb{Z}}=\mathbb{V}$. Let $\mathbb{W}=\bigwedge^{3}_{\mathcal{O}}R^{1}\pi_{*}{\rm\mathbb{Z}}$ (the third exterior product of $R^{1}\pi_{*}{\rm\mathbb{Z}}$, viewed as a sheaf of $\mathcal{O}$-modules), a ${\rm\mathbb{Z}}$-VHS with an $\mathcal{O}$-action. Lemma 4.6 shows that $\mathbb{W}\otimes{\rm\mathbb{Q}}\simeq\mathbb{V}^{\mu_{3}}\otimes{\rm\mathbb{Q}}$, so $\mathrm{J}(\mathbb{W})$ and $\mathrm{J}(\mathbb{V}^{\mu_{3}})$ are isogenous.
To analyze $\mathrm{J}(\mathbb{W})$, we analyze the ${\rm\mathbb{Z}}$-VHS $\mathbb{W}(1)$ more closely. It has an $\mathcal{O}$-action by construction, Lemma 4.6 shows that it has constant rank $2$, and Lemma 4.5 shows that it has type $(1,0)+(0,1)$. Since $\mathcal{O}$ has class number $1$, there exists a unique ${\rm\mathbb{Z}}$-Hodge structure with these properties, hence $\mathbb{W}_{f}(1)\simeq\operatorname{H}^{1}(E;{\rm\mathbb{Z}})$ for every $f\in S({\rm\mathbb{C}})$, where $E$ is the elliptic curve with Weierstrass equation $y^{2}=x^{3}+1$.
Therefore $\mathrm{J}(\mathbb{W})\rightarrow S({\rm\mathbb{C}})$ is an isotrivial family of elliptic curves: analytically locally on $S({\rm\mathbb{C}})$, it is isomorphic to $E({\rm\mathbb{C}})\times S({\rm\mathbb{C}})\rightarrow S({\rm\mathbb{C}})$. It follows that the sheaf of local isomorphisms between $\mathrm{J}(\mathbb{W})$ and $E({\rm\mathbb{C}})\times S({\rm\mathbb{C}})$ is an ${\rm Aut}(E)$-torsor in the analytic topology on $S({\rm\mathbb{C}})$. Since ${\rm Aut}(E)\simeq\mu_{6}$ is finite, this torsor is the analytification of an étale $\mu_{6}$-torsor on $S$.
To analyze étale $\mu_{6}$-torsors on $S$, consider the following exact sequence induced by the Kummer exact sequence in étale cohomology:
${\rm\mathbb{G}}_{m}(S)\xrightarrow{(-)^{6}}{\rm\mathbb{G}}_{m}(S)\rightarrow\operatorname{H}^{1}_{\mathrm{et}}(S,\mu_{6})\rightarrow{\rm Pic}(S)[6]\rightarrow 0.$
The Picard group ${\rm Pic}(S)$ vanishes, being a quotient of ${\rm Pic}({\rm\mathbb{A}}^{3}_{{\rm\mathbb{C}}})$, hence the outer term in the sequence vanishes. We claim that ${\rm\mathbb{G}}_{m}(S)=\operatorname{H}^{0}(S,\mathcal{O}_{S})^{\times}=\{c\cdot{\rm disc}^{n}\mid c\in{\rm\mathbb{C}}^{\times},\,n\in{\rm\mathbb{Z}}\}$. Indeed, every $c\cdot{\rm disc}^{n}$ is clearly a unit in $\operatorname{H}^{0}(S,\mathcal{O}_{S})$. Conversely, given a unit $f\in\operatorname{H}^{0}(S,\mathcal{O}_{S})$, seen as a rational function on ${\rm\mathbb{A}}^{3}_{{\rm\mathbb{C}}}$, its divisor $\mathrm{div}(f)$ of zeros and poles must be supported on the zero locus of ${\rm disc}$. Since ${\rm disc}\in{\rm\mathbb{C}}[a,b,c]$ is irreducible, $\mathrm{div}(f)=n\cdot[\{{\rm disc}=0\}]$ for some $n\in{\rm\mathbb{Z}}$. Then $\mathrm{div}(f/{\rm disc}^{n})=0$ as a rational function on ${\rm\mathbb{A}}^{3}_{{\rm\mathbb{C}}}$, hence $f/{\rm disc}^{n}$ is a unit in ${\rm\mathbb{C}}[a,b,c]$, hence $f/{\rm disc}^{n}\in{\rm\mathbb{C}}^{\times}$, proving the claim. We conclude that the group $\operatorname{H}^{1}_{\mathrm{et}}(S,\mu_{6})$ classifying $\mu_{6}$-torsors is generated by the image of ${\rm disc}$.
Let $\mathcal{E}_{i}\rightarrow S$ be the relative elliptic curve with equation $y^{2}=x^{3}+{\rm disc}(f)^{i}$. The previous paragraph shows that $\mathrm{J}(\mathbb{W})$ is isomorphic to $\mathcal{E}_{i}({\rm\mathbb{C}})$ for some $i\in\{0,1,2,3,4,5\}$. We show that $i=1$, using our previous results on bielliptic Picard curves in [27]. Let $T\subset S$ be the closed subscheme where $b=0$, parametrizing even quartic polynomials $f=x^{4}+ax^{2}+c$. For $f\in T({\rm\mathbb{C}})$, the $\mu_{3}$-action on $\mathcal{C}_{f}$ extends to a $\mu_{6}$-action. A calculation shows ${\rm disc}|_{T}=16c(-a^{2}+4c)^{2}$. Applying the singular cohomology realization functor to [27, Theorem 5.1], the relative complex torus $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})\rightarrow T({\rm\mathbb{C}})$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}\rightarrow T({\rm\mathbb{C}})$. Since $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{6}})$ is a subtorus of $\mathrm{J}((\mathbb{V}|_{T({\rm\mathbb{C}})})^{\mu_{3}})$ of the same dimension, they must be equal, hence $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ over $T({\rm\mathbb{C}})$. On the other hand, let $i\in\{0,1,2,3,4,5\}$ be such that $\mathrm{J}(\mathbb{W})\simeq\mathcal{E}_{i}({\rm\mathbb{C}})$. Then $\mathrm{J}(\mathbb{W})|_{T({\rm\mathbb{C}})}\simeq\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$, hence $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ is isogenous to $\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$.
We show that the latter can happen only if $i=1$. Indeed, let $\varphi\colon\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}\rightarrow\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ be an isogeny. Since the domain and target of $\varphi$ are isotrivial relative elliptic curves with $\mathcal{O}$-multiplication, $\varphi$ factors as $\psi\circ\gamma$, where $\gamma$ is an endomorphism of $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ and $\psi$ an isomorphism. Therefore $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ and $\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ are isomorphic. Since the monodromy representations of $\mathcal{E}_{1}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ and $\mathcal{E}_{i}({\rm\mathbb{C}})|_{T({\rm\mathbb{C}})}$ are non-isomorphic when $i\neq 1$, we conclude that $i=1$ and that $\mathrm{J}(\mathbb{W})$ is isogenous to $\mathcal{E}_{1}({\rm\mathbb{C}})$ over $S({\rm\mathbb{C}})$.
In summary, we have shown that $\mathrm{J}(\mathbb{V}^{\mu_{3}})$, $\mathrm{J}(\mathbb{W})$ and $\mathcal{E}_{1}({\rm\mathbb{C}})$ are isogenous. Since $\mathcal{E}_{1}({\rm\mathbb{C}})$ is isomorphic to $\mathcal{E}({\rm\mathbb{C}})$, we conclude the proof. ∎
Proposition 4.15.
The group of (algebraic) sections of $\mathcal{E}\rightarrow S$ is free of rank $1$ over $\mathcal{O}={\rm\mathbb{Z}}[\omega]$, and contains the $\mathcal{O}$-span of $P:=(a^{2}+12c,72ac-2a^{3}-27b^{2})$ as a finite index subgroup.
Proof.
Given $s=(a,b,c)\in S({\rm\mathbb{C}})$, consider the closed subscheme $T=\{(at,bt,ct^{2})\colon t\in{\rm\mathbb{A}}^{1}_{{\rm\mathbb{C}}}\}\cap S$ of $S$ and the restriction $\pi\colon\mathcal{E}|_{T}\rightarrow T$. The variety $\mathcal{E}|_{T}$ is an open subscheme of an elliptic surface with Weierstrass equation $y^{2}=x^{3}+t^{4}g(t)$, where $g(t)\in{\rm\mathbb{C}}[t]$ has degree $\leq 2$, using the formula (4.1). There exists a dense open $U\subset S$ such that for all $s\in U({\rm\mathbb{C}})$, $g(t)$ has two distinct nonzero roots. Let $s\in U({\rm\mathbb{C}})$. Using Tate’s algorithm, we see that the elliptic surface has a singular fiber above $t=0$ with Kodaira type $\text{\rm IV}^{*}$, singular fibers above the roots of $g(t)$ with Kodaira type II, and is smooth above $t=\infty$. The presence of a fiber of type II implies that $\mathcal{E}|_{T}\rightarrow T$ has no torsion sections [43, Lemma 7.8]. Moreover the Shioda–Tate formula [43, Theorem 6.3, Proposition 6.6 and §8.8] shows that the group of sections of $\mathcal{E}|_{T}\rightarrow T$ is free of rank $2$. Since $\mathcal{E}|_{T}$ receives an $\mathcal{O}$-action, its group of sections is free of rank $1$ over $\mathcal{O}$. We will now show that these facts can be used to prove the claims of the proposition by varying $s$ in $U({\rm\mathbb{C}})$.
We first show that $\mathcal{E}(S)$ is torsion-free. Suppose $Q\in\mathcal{E}(S)$ is a torsion section. Then $N\cdot Q=0$ for some $N\geq 1$. Hence $N\cdot Q|_{T}=0$ for every $s\in U({\rm\mathbb{C}})$. Since the group of sections of $\mathcal{E}|_{T}\rightarrow T$ is torsion-free, $Q|_{T}=0$ for every $s\in U({\rm\mathbb{C}})$. Hence $Q|_{U}=0$. Since $U$ is dense in $S$, we must have $Q=0$, as desired.
The fact that $P$ defines a section of $\mathcal{E}\rightarrow S$ can be verified by direct computation (see §4.1). Let $L={\rm\mathbb{Z}}\langle P,\omega\cdot P\rangle$ be the $\mathcal{O}$-span of $P$, which is a subgroup of the group of sections $\mathcal{E}(S)$ of $\mathcal{E}\rightarrow S$. The previous paragraph shows that $L$ is free of rank $1$ over $\mathcal{O}$. It remains to show that it has finite index in $\mathcal{E}(S)$. Suppose for the sake of contradiction that there exists a third section $Q\in\mathcal{E}(S)$ which is not in $L\otimes{\rm\mathbb{Q}}$. Then the locus $S_{\text{dep}}$ of $s\in S({\rm\mathbb{C}})$ where $P_{s},\omega\cdot P_{s}$ and $Q_{s}$ are linearly dependent is a countable union of closed proper algebraic subvarieties. For every $s\in U({\rm\mathbb{C}})$, the group of sections of $\mathcal{E}|_{T}\rightarrow T$ is free of rank $1$ over $\mathcal{O}$. There exists a possibly smaller dense open $V\subset U$ such that if $s\in V({\rm\mathbb{C}})$ then $P_{s}$ is nonzero, hence $\langle P_{s},\omega\cdot P_{s}\rangle$ is a finite index subgroup of $\mathcal{E}(T)$. Therefore $Q_{t},P_{1,t},P_{2,t}$ are linearly dependent for all $s\in V({\rm\mathbb{C}})$. Since $V({\rm\mathbb{C}})\setminus S_{\text{dep}}$ is nonempty, we obtain a contradiction. ∎
4.7. Proof of the vanishing criterion in the Chow group
Proof of Theorem 4.2.
We may assume (using Lemma 2.1) that $k={\rm\mathbb{C}}$. Recall that $3\cdot{\rm AJ}(\kappa_{\mathcal{C}})$ defines a holomorphic section of $\mathrm{J}(\mathbb{V}^{\mu_{3}})\rightarrow S({\rm\mathbb{C}})$. Choose an isogeny of complex tori $\mathrm{J}(\mathbb{V}^{\mu_{3}})\rightarrow\mathcal{E}({\rm\mathbb{C}})$ using Proposition 4.14 and let $\sigma$ be the image of $3\cdot{\rm AJ}(\kappa_{\mathcal{C}})$ under this isogeny. This is a holomorphic section of $\mathcal{E}({\rm\mathbb{C}})\rightarrow S({\rm\mathbb{C}})$.
We claim that $\sigma$ is not a torsion section. If it were torsion, then ${\rm AJ}(\kappa_{\mathcal{C}})$ would be a torsion section of $\mathrm{J}(\mathbb{V}^{\mu_{3}})$, hence, by Proposition 4.12, $\kappa_{f}$ would be torsion for every $f\in S({\rm\mathbb{C}})$. This is not the case, since $\kappa_{f}$ is of infinite order if $f=x^{4}+x^{2}+1$ by [27, Corollary 2.9]. We conclude that $\sigma$ is not a torsion section.
Next we claim that $\sigma$ is the analytification of an algebraic section of $\mathcal{E}\rightarrow S$. Let $N\geq 1$ be an integer such that $N\cdot\kappa_{f}$ is algebraically trivial for every Picard curve $f\in S({\rm\mathbb{C}})$ (such an integer exists by Theorem 4.1). By the algebraicity of the Abel–Jacobi map for algebraically trivial cycles ([1, Theorem 1]), there exists a relative algebraic subtorus $\mathrm{J}_{a}(\mathbb{V})\subset\mathrm{J}(\mathbb{V})$ with the following property: the section $\mathrm{AJ}(3N\cdot\kappa_{\mathcal{C}})$ lands in $\mathrm{J}_{a}(\mathbb{V})$ and the corresponding holomorphic map $S({\rm\mathbb{C}})\rightarrow\mathrm{J}_{a}(\mathbb{V})$ is algebraic. On the other hand, ${\rm AJ}(3N\cdot\kappa_{\mathcal{C}})$ also lands in $\mathrm{J}(\mathbb{V}^{\mu_{3}})$, is not a torsion section by the previous paragraph, and $\mathrm{J}(\mathbb{V}^{\mu_{3}})$ has relative dimension $1$ over $S({\rm\mathbb{C}})$. Therefore $\mathrm{J}(\mathbb{V}^{\mu_{3}})$ is the smallest relative subtorus of $\mathrm{J}(\mathbb{V})$ containing the image of ${\rm AJ}(3N\cdot\kappa_{\mathcal{C}})$. Hence $\mathrm{J}(\mathbb{V}^{\mu_{3}})\subset\mathrm{J}_{a}(\mathbb{V})$. We conclude that ${\rm AJ}(3N\cdot\kappa_{\mathcal{C}})\colon S({\rm\mathbb{C}})\rightarrow\mathrm{J}(\mathbb{V}^{\mu_{3}})$ is algebraic, so ${\rm AJ}(3\cdot\kappa_{\mathcal{C}})$ is algebraic, so $\sigma$ is algebraic.
Since $\sigma$ is an algebraic section of $\mathcal{E}\rightarrow S$, Proposition 4.15 shows that there exists an integer $M\geq 1$ and an element $\gamma\in\mathcal{O}$ such that $M\cdot\sigma=\gamma\cdot P$. Since $\sigma$ is not a torsion section, $\gamma\neq 0$.
Putting everything together, we have for $f\in S({\rm\mathbb{C}})$: $\kappa_{f}$ is torsion if and only if ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$ torsion (by Proposition 4.12), if and only if ${\rm AJ}(\kappa_{\mathcal{C}})_{f}\in\mathrm{J}(\mathbb{V}^{\mu_{3}})_{f}$ torsion, if and only if $\sigma_{f}\in\mathcal{E}_{f}({\rm\mathbb{C}})$ torsion, if and only if $P_{f}$ torsion. Tracing through the equivalences, the quotient of the torsion orders $\operatorname*{ord}(\kappa_{f})/\operatorname*{ord}(P_{f})$ (if defined) takes only finitely many values as $f$ ranges in $S({\rm\mathbb{C}})$. ∎
4.8. Families of Picard curves with torsion Ceresa cycle of arbitrarily large order
Theorem 4.2 shows that there are infinitely many plane quartic curves over ${\rm\mathbb{Q}}$ with torsion Ceresa cycle, since we may take $a=c=-12$. In fact, we can find explicit families of torsion Ceresa cycles of arbitrarily large order by analyzing the function $f\mapsto(I(f),J(f))$.
Let $C_{f}$ be a Picard curve defined by the polynomial $f(x)=x^{4}+ax^{2}+bx+c$ over a field $k$ with homogenization $F(x,z)=x^{4}+ax^{2}z^{2}+bxz^{3}+cz^{4}$. Associated to $C_{f}$ is the genus $1$ curve $D_{f}$ with equation $y^{2}=F(x,z)$. Viewing $D_{f}$ as an elliptic curve with origin $(1:1:0)$, it is isomorphic to the elliptic curve $y^{2}=x^{3}-I(f)x/3-J(f)/27$. The point $(1:-1:0)$ is sent to $(-2a/3,b)$ in this new model. Conversely, given a nonzero $k$-point $(\alpha,\beta)$ on the elliptic curve $E_{I,J}\colon y^{2}=x^{3}-Ix/3-J/27$, the quartic
(4.3) $f(x)=x^{4}-3\alpha x^{2}/2+\beta x+(I/12-3\alpha^{2}/16)$
has the property that $(I(f),J(f))=(I,J)$. It follows that if $(I,J)\in k$ is such that $4J^{3}-J^{2}\neq 0$, then $f\mapsto(-2a/3,b)$ induces a bijection between the set of polynomials $f=x^{4}+ax^{2}+bx+c\in k[x]$ satisfying $(I(f),J(f))=(I,J)$, and the set of nonzero elements of $E_{I,J}(k)$. Using this description, we can cook up explicit 1-parameter families of torsion Ceresa cycles of arbitrarily large order, as in the following corollary of Theorem 4.2:
Corollary 4.16.
For every integer $N$, there exists a number field $k=k(N)$ such that there exists infinitely many $\bar{\rm\mathbb{Q}}$-isomorphism classes Picard curves $C_{f}$ defined over $k$ with Ceresa cycle $\kappa_{f}$ of order at least $N$ in ${\rm CH}^{2}(J_{f,\bar{k}})$.
Proof.
Let $M$ be an integer satisfying the conclusion of Theorem 4.2. Fix a nonzero $D\in{\rm\mathbb{Q}}$ and let $(I,J)\in\bar{{\rm\mathbb{Q}}}^{2}$ be a torsion point of order $M\cdot N$ on the elliptic curve $y^{2}=4x^{3}+D$. Let $k$ be the field of definition of $(I,J)$. For $t\in k$, define $\lambda(t)=t^{3}-It/3-J/27$, $(\alpha_{t},\beta_{t})=(t\lambda(t),\lambda(t)^{2})$ and the quartic $f_{t}(x)=x^{4}-3\alpha_{t}x^{2}/2+\beta_{t}x+(I/12-3\alpha_{t}^{2}/16)$ using Formula (4.3). Since $(\alpha_{t},\beta_{t})$ defines a point on the elliptic curve $E_{\lambda(t)^{2}I,\lambda(t)^{3}J}$, the quartic $f$ has the property that $(I({f_{t}}),J({f_{t}}))=(\lambda(t)^{2}I,\lambda(t)^{3}J)$. If $t\in k$ is such that $\lambda(t)\neq 0$, then $(\lambda(t)^{2}I,\lambda(t)^{3}J)$ is an $N$-torsion point on the elliptic curve $y^{2}=4x^{3}+\lambda_{t}^{6}D$. By Theorem 4.2, $\kappa_{f_{t}}$ is torsion in ${\rm CH}^{2}(J_{f_{t},\bar{k}})$ and its torsion order $\operatorname*{ord}(\kappa_{f_{t}})$ is divisible by $N$. When varying over $t\in k$ such that $\lambda(t)\neq 0$, the Picard curves $C_{f_{t}}$ cover infinitely many $\bar{\rm\mathbb{Q}}$-isomorphism classes, using [13, Lemma 1.21(b)] and the fact that $\beta_{t}^{2}/\alpha_{t}^{3}=\lambda(t)/t^{3}$ is not a constant function of $t$. ∎
Conversely, for any number field $k$, the order of a torsion Ceresa cycle $\kappa(C_{f})$, with $C_{f}$ defined over $k$, is bounded, with the bound depending only on the degree of $k$ over ${\rm\mathbb{Q}}$:
Corollary 4.17.
For every $d\geq 1$, there exists $N=N(d)\geq 1$ such that for every Picard curve $C_{f}$ over a number field $k$ of degree $d$, the order of $\kappa_{f}$ in ${\rm CH}_{1}(J_{f,\bar{k}})$ is either infinite or less than $N$.
Proof.
This follows from Theorem 4.2 and the uniform bound (depending only on $d$) on the order of a $k$-rational torsion point on an elliptic curve $y^{2}=x^{3}+D$ over any number field $k$ of degree $d$. ∎
5. Automorphism strata and Ceresa vanishing loci in genus $3$
Fix $g\geq 3$ and let $\mathcal{M}_{g}$ be the (coarse) moduli space of genus-$g$ curves, seen as a variety over ${\rm\mathbb{Q}}$. Let $V_{g}^{\mathrm{rat}}\subset\mathcal{M}_{g}$ be the subset of curves $[C]$ for which $\kappa(C)$ vanishes in ${\rm CH}_{1}(\operatorname{Jac}(C))_{{\rm\mathbb{Q}}}$, in the notation of §2.7. Since the vanishing of $\kappa(C)$ only depends on the geometric isomorphism class of $C$ (by Lemma 2.1), this locus is well defined. Similarly define the locus $V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ where $\bar{\kappa}(C)\in{\rm Gr}_{1}(\operatorname{Jac}(C))_{{\rm\mathbb{Q}}}$ vanishes.
Lemma 5.1.
The subsets $V_{g}^{\mathrm{rat}},V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$ are countable unions of proper closed algebraic subvarieties.
Proof.
Let ~$\widetilde{\mathcal{M}}_{g}$ be the fine moduli space parametrizing genus-$g$ curves $C$ with full symplectic level-$5$ structure and a degree-$1$ divisor class $e\in{\rm CH}_{0}(C)$ such that $(2g-2)e$ is canonical. Considering the universal curve over it together with its degree-$1$ divisor, we can define a universal Ceresa cycle on the universal Jacobian over ~$\widetilde{\mathcal{M}}_{g}$; Lemma 2.4 then implies that the locus in ~$\widetilde{\mathcal{M}}_{g}$ where this Ceresa cycle vanishes (with ${\rm\mathbb{Q}}$-coefficients) is a countable union of closed algebraic subvarieties. Since the forgetful map ~$\widetilde{\mathcal{M}}_{g}\rightarrow\mathcal{M}_{g}$ is proper, the same is true for the image of this locus, which is exactly $V_{g}^{\mathrm{rat}}$. The proof for $V_{g}^{\mathrm{alg}}$ is identical. ∎
These vanishing loci have the following basic properties: $V_{g}^{\mathrm{rat}}\subset V_{g}^{\mathrm{alg}}\subset\mathcal{M}_{g}$; the hyperelliptic locus is contained in $V_{g}^{\mathrm{rat}}$; and $V_{g}^{\mathrm{alg}}\neq\mathcal{M}_{g}$ by Ceresa’s famous result [14]. It would be interesting to obtain further information about the components of $V_{g}^{\mathrm{rat}}$ and $V_{g}^{\mathrm{alg}}$. We end our paper by determining the automorphism group strata in $\mathcal{M}_{3}$ that are contained in $V_{3}^{\mathrm{rat}}$ or $V_{3}^{\mathrm{alg}}$.
So let $g=3$ and consider the open subscheme $\mathcal{M}_{3}^{\mathrm{nh}}\subset\mathcal{M}_{3}$ of non-hyperelliptic curves. There is a stratification $\mathcal{M}_{3}^{\mathrm{nh}}=\sqcup X_{G}$ into locally closed subvarieties such that a non-hyperelliptic curve $C$ over ${\rm\mathbb{C}}$ belongs to $X_{G}$ if and only if ${\rm Aut}(C)\simeq G$. It turns out that $X_{G}$ is irreducible and the closure of $X_{G}$ is a union of other strata. We refer to [31, §2.2] and references therein for a complete description of the loci $X_{G}$ and the closure relations between them. We reproduce here a diagram capturing these closure relations:
[…]
For $n\in\{16,48,96\}$, the symbol $G_{n}$ the group of order $n$ and GAP label $(16,13),(48,33),$ and $(96,64)$ respectively [21]. See [31, Table 2] for models for a generic plane quartic in $X_{G}$. We make explicit the strata that are relevant for us: $\overline{X_{C_{3}}}$ is the locus of Picard curves studied in §4; $\overline{X_{C_{6}}}$ is the locus of bielliptic Picard curves studied in [28]; and the zero-dimensional strata each consist of a single automorphism-maximal curve with equation
$\begin{cases}y^{3}z=x^{4}+xz^{3}&\text{ if }G=C_{9},\\ y^{3}z=x^{4}+z^{4}&\text{ if }G=G_{48,}\\ x^{4}+y^{4}+z^{4}=0&\text{ if }G=G_{96},\\ x^{3}y+y^{3}z+z^{3}x&\text{ if }G={\rm GL}_{3}({\rm\mathbb{F}}_{2}).\end{cases}$
If $C$ is a non-hyperelliptic genus $3$ curve over a field $k$, we say $C$ is a generic curve for $X_{G}$ if the classifying map ${\rm Spec}(k)\rightarrow\mathcal{M}_{3}$ maps to the generic point of $X_{G}$. If $X$ is an integral variety over ${\rm\mathbb{C}}$, we say a property hold for a very general $x\in X({\rm\mathbb{C}})$ if it holds true outside a countable union of proper closed subvarieties of $X$.
Lemma 5.2.
For a group $G$ in the above diagram, the following are equivalent:
  1. (1)
    $\kappa(C)\neq 0$ for some generic curve $C$ for $X_{G}$.
  2. (2)
    $\kappa(C)\neq 0$ for a very general $C$ in $X_{G}({\rm\mathbb{C}})$.
  3. (3)
    $\kappa(C)\neq 0$ for some $C$ in $X_{G}$.
  4. (4)
    $X_{G}\not\subset V_{3}^{\mathrm{rat}}$.
  5. (5)
    $\overline{X_{G}}\not\subset V_{3}^{\mathrm{rat}}$.
Moreover, the analogous equivalences hold for $\bar{\kappa}(C)\in{\rm Gr}_{1}(J)_{\rm\mathbb{Q}}$ and $V_{3}^{\mathrm{alg}}$.
Proof.
Follows immediately from Lemma 5.1. ∎
We end with the proof of Theorem D, which we restate for convenience:
Proposition 5.3.
Let $G$ be a group in the diagram. Then
  1. (1)
    $X_{G}\subset V_{3}^{\mathrm{alg}}$ if and only if $G=C_{3},C_{6},C_{9},G_{48}$.
  2. (2)
    $X_{G}\subset V_{3}^{\mathrm{rat}}$ if and only if $G=C_{9},G_{48}$;
Proof.
  1. (1)
    The equivalence between $(4)$ and $(5)$ of Lemma 5.2 implies that if $X_{H}\subset\overline{X_{G}}$ and $X_{G}\subset V_{3}^{\mathrm{rat}}$, then $X_{H}\subset V_{3}^{\mathrm{rat}}$; we will use this observation in the remainder of the proof.
    Our analysis of Picard curves (Theorem C) shows that $X_{C_{3}}\subset V_{3}^{\mathrm{rat}}$, so $X_{G}\subset V_{3}^{\mathrm{rat}}$ for $G=C_{6},C_{9}$ and $G_{48}$ as well. On the other hand, the Ceresa cycle of the Fermat quartic and Klein quartic are known to be of infinite order in the Griffiths group; see [11, Theorem (4.1)] for the former and [24, §4] for the latter. By our observation, this means that $X_{G}\not\subset V_{3}^{\mathrm{rat}}$ for every stratum whose closure contains one of these curves. Since every stratum not contained in $\overline{X_{C_{3}}}$ has this property, we conclude the proof.
  2. (2)
    Since $V_{3}^{\mathrm{rat}}\subset V_{3}^{\mathrm{alg}}$, Part $(1)$ implies that $X_{G}\subset V_{3}^{\mathrm{rat}}$ only if $G=C_{3},C_{6},C_{9}$ or $G_{48}$. The criterion of Theorem A applies to the curves in $X_{C_{9}}$ and $X_{G_{48}}$ (see Example 3.1), so $V_{3}^{\mathrm{rat}}$ contains these strata. On the other hand, there exist curves in $X_{C_{6}}$ with nonvanishing $\kappa(C)$, by [27, Corollary 1.2]. So $X_{C_{6}}\not\subset V_{3}^{\mathrm{rat}}$ and $X_{C_{3}}\not\subset V_{3}^{\mathrm{rat}}$.
Remark 5.4.
Not every irreducible component of $V_{3}^{\mathrm{rat}}$ is of the form $\overline{X_{G}}$ for some $G$. The proof of Corollary 4.16 shows that $V_{3}^{\mathrm{rat}}\cap\overline{X_{C_{3}}}$ is a union of countably many (open, possibly singular) rational curves.